### v2

# Low-Dimensional Adaptation of Rectified Flow: A Diffusion and Stochastic Localization Perspective

Abstract
In recent years, Rectified flow (RF) has gained considerable popularity largely due to its generation efficiency and state-of-the-art performance. In this paper, we investigate how well RF automatically adapts to the intrinsic low dimensionality of the support of the target distribution to accelerate sampling. We show that, using a carefully designed choice of the time-discretization scheme and with sufficiently accurate drift estimates, the RF sampler enjoys an iteration complexity of order $O(k/\varepsilon)$ (up to log factors), where $\varepsilon$ is the precision in total variation distance and $k$ is the intrinsic dimension of the target distribution. In addition, we show that the denoising diffusion probabilistic model (DDPM) procedure is equivalent to a stochastic version of RF by establishing a novel connection between these processes and stochastic localization. Building on this connection, we further design a stochastic RF sampler that also adapts to the low-dimensionality of the target distribution under mild requirements on the accuracy of the drift estimates, and also with a specific time schedule. We illustrate the efficacy of newly designed time-discretization schedules with simulations on the synthetic data and text-to-image (T2I) data experiments.
Machine Learning, ICML
Department of Statistics and Data Science
University of Texas at Austin, TX, USA
1 Introduction
Recently, Rectified flow or Flow matching (Liu, 2022; Lipman et al., 2022) has gained much attention due to its state-of-the-art generative performance across various modalities including image (Liu et al., 2024; Yan et al., 2024), audio (Wang et al., 2024; Liu et al., 2025a), and video (Chen et al., 2025; Liu et al., 2025b). Its effectiveness arises from a deterministic ordinary differential equation (ODE) based sampler and the ability of reflow to learn nearly straight transport paths, resulting in faster sampling. This has naturally sparked theoretical interests (Bansal et al., 2025; Guan et al., 2025; Zhou and Liu, 2025) to understand the generation quality of RF from a convergence viewpoint under a suitable metric. However, these recent results suffer from curse of dimensionality. Therefore, they fail to explain the impressive performance of RF for high-dimensional data, e.g., image data, even though they reside in a low-dimensional space (Ansuini et al., 2019; Pope et al., 2021). In contrast, recent works (Potaptchik et al., 2024; Azangulov et al., 2024; Huang et al., 2024; Li and Yan, 2024; Liang et al., 2025) have established low-dimensional adaptation of DDPM and denoising diffusion implicit models (DDIM) during sampling. This explains the accelerated and high-quality generative capabilities of DDPM and DDIM in practice, even for high-dimensional data. To the best of our knowledge such results are not currently present for RF models. To this end, we make three key contributions in this paper:
  1. 1.
    We design a novel U-shaped time-discretization that provably allows RF to quickly achieve improved generation quality in the first reflow process by adapting to the intrinsic low-dimensionality of the target distribution. This aligns with the recent work by Lee et al. (2024), which has empirically shown that a certain U-shaped time-discretization improves the generative quality of the first reflow model. In particular, the RF sampler enjoys $O(k/\varepsilon)$ (upto log factors) iteration complexity, where $\varepsilon$ is the precision of the generated samples in total variation distance and $k$ is the intrinsic dimension of the target distribution. This is particularly useful because it can eliminate repeated reflow steps, which are computationally costly and often cause model collapse (Zhu et al., 2024; Kim et al., 2025).
  2. 2.
    Using a novel stochastic localization (Eldan, 2013, 2020) perspective, we demonstrate that DDPM is equivalent to the stochastic version of RF (which we refer to as STOC-RF) under a proper time change. Stochastic localization has been recently shown to be provide a unifying and effective framework for representing and analyzing diffusion-based sampling processes (Benton et al., 2023; Montanari, 2023; El Alaoui and Montanari, 2022). This equivalence is convenient as it allows us to view STOC-RF through the lens of well-understood DDPM processes, which are known to adapt to the low-dimensionality of the target distribution. In fact, our analysis shows that stochastic localization can be leveraged to provide a unified and convenient representation of a variety of sampling methods.
  3. 3.
    Using the connection between DDPM and STOC-RF, we further design a stochastic RF sampler that provably also adapts to the low-dimensional structure of the target distributions under less stringent assumptions on the learned model. We illustrate the superior performance of this sampler with various simulations. This also addresses one of the themes discussed in the ICML tutorial talk by Liu (2025).
Notation. Throughout the paper, with some abuse of notations, we will identify a probability distribution with its Lebesgue density, which we will assume exists unless otherwise noted. For two distributions $P$ and $Q$ on $\mathbb{R}^{d}$ with Lebesgue densities $p$ and $q$ respectively, their total variation distance is denoted by ${\sf TV}(p,q):=\sup_{A}|{P(A)-Q(A)}|=1/2\int_{\mathbb{R}^{d}}|p(x)-q(x)|dx$ , where the supremum is over all Borel subsets of $\mathbb{R}^{d}$.
2 Background and Preliminaries
2.1 Rectified Flow
In this section, we briefly introduce the basics of Rectified flow, a generative model that transitions between two distributions $p_{0}$ and $p_{1}$ by solving ODEs. We refer the reader to Liu (2024, 2022) for an exhaustive treatment. We let $p_{1}:=\text{Law}(X_{1})$ be the target data distribution with its support $\mathcal{X}\subseteq\mathbb{R}^{d}$, the data generating distribution from which we would like to draw samples. The linear-interpolation process is defined by by
$X_{t}=tX_{1}+(1-t)X_{0},\quad 0\leq t\leq 1$ (1)
where $p_{0}$ is the standard Gaussian distribution. We will assume an independent initial coupling, i.e., $(X_{0},X_{1})\sim p_{0}\otimes p_{1}$ and let $p_{t}=\text{Law}(X_{t})$, for any $t\in[0,1]$. The RF procedure is a deterministic sampler arising from an ODE with drift (or velocity) function $v:\mathbb{R}^{d}\times[0,1]\rightarrow\mathbb{R}^{d}$ defined as the solution to the optimization problem
$v$ $={\arg\min}_{f}\int\mathbb{E}{\|X_{1}-X_{0}-f(X_{t},t)\|_{2}^{2}}dt,$ (2)
where the minimization is over all functions $f:\mathbb{R}^{d}\times[0,1]\rightarrow\mathbb{R}^{d}$. For each $t$, The above objective in (2) is minimized at $(x,t)\mapsto v_{t}(x):=v(x,t)=\mathbb{E}[X_{1}-X_{0}\mid X_{t}=x].$ Under appropriate regularity conditions (investigated in Bansal et al., 2025; Mena et al., 2025), the ODE
$dZ_{t}=v_{t}(Z_{t})\ dt,\quad Z_{0}\sim N(0,I_{d})$ (3)
satisfies the marginal preserving property, i.e., $\text{Law}(Z_{t})=\text{Law}(X_{t})=p_{t}$, owing to the Fokker-Planck equation
$\frac{\partial p_{t}}{\partial t}+\nabla.(v_{t}p_{t})=0.$ (4)
Hence, ODE (3) can be used for sampling.
To implement this procedure, one faces two key challenges. First, given a sample from the target distribution $p_{1}$, one has to estimate the drift function, which is naturally done by empirically minimizing the objective in (2) over a large and expressive function class (e.g., U-Net). Secondly, for sampling, one must rely on a time-discretization scheme of the ODE (3). One such scheme is the Eulerian update rule, which, for a given discretization of the time course in $N$ intervals, with $0=t_{0}<t_{1}<\dots<t_{N}=1$, evaluates
$Y_{t_{i+1}}=Y_{t_{i}}+\eta_{i}\widehat{v}_{t_{i}}(Y_{t_{i}});\;\eta_{i}:=t_{i+1}-t_{i},0\leq i\leq N-1,$ (5)
where $Y_{0}\sim N(0,I_{d})$ and $\widehat{v}_{t}(\cdot)$ is the estimated velocity at time $t$. Let $\widehat{p}_{1}:=\text{Law}(Y_{1})$ denote the distribution of the final step. Other types of time discretized schemes of the ODE (3) can be chosen for the sampling step, and the resulting discretization errors impact the sampling fidelity differently. This has been the focus of a recent line of works (Benton et al., 2023; Chen et al., 2023; Li et al., 2024a, b; Huang et al., 2024; Liang et al., 2025) aimed at identifying time-discretized versions of DDPM and DDIM samplers to achieve tight convergence rates.
The velocity function.
Before proceeding, we illustrate the connection between the velocity $v_{t}(x)$ and the score function $s_{t}(x):=\nabla\log p_{t}(x)$. Recall that $X_{0}\sim N(0,I_{d})$ and independent of $X_{1}\sim p_{1}$. Therefore, Tweedie’s formula (Efron, 2011; Meng et al., 2021) immediately yields $v_{t}(x)=\frac{x}{t}+\frac{(1-t)}{t}s_{t}(x)$. Thus, estimating $v_{t}$ is essentially equivalent to estimating the score $s_{t}$ for $t\in(0,1)$.
2.2 Stochastic rectified flow (STOC-RF)
Since the work of DDIM (Song et al., 2021a) and DDPM (Song et al., 2021b), it is well known that SDEs can be converted to ODEs to obtain deterministic samplers. Conversely, it is also possible to convert ODEs to its equivalent SDE form to obtain stochastic samplers. In fact, prior works (Albergo et al., 2023; Xue et al., 2024; Hu et al., 2025a) have shown that stochastic samplers typically enjoy better generation quality. One reason behind the inferior performance of deterministic samplers is the possibility of error accumulation during the sampling step along the discrete trajectories stemming from inaccurate velocity estimates. In other words, if $Y_{t_{i}}$ differs significantly from $Z_{t_{i}}$ for some $t_{i}$, then the Euler step (5) accumulates error due to inaccurate evaluation of $\widehat{v}_{t_{i}}(Y_{t_{i}})$. Therefore, this can further reinforce the deviation in evaluating $Y_{t_{i+1}}$. A way to mitigate the compounding error from successive Euler steps is to introduce a stochastic Langevin correction at each step. Empirically, Hu et al. (2025b) showed that the resulting stochastic RF samplers enjoy better text rendering quality over deterministic samplers. We further illustrate this point with Figure 1, showing that the ODE-based RF trajectories appear to generate more outliers compared to the SDE-based STOC-RF sampler.
Figure 1: Trajectories of RF and STOC-RF samplers for a mixture of 2-Gaussian target distribution.
Following e.g. Liu (2024, Chapter 5), the stochastic version of RF is specified by the SDE
$d\tilde{Z}_{t}=v_{t}(\tilde{Z}_{t})dt+\underbrace{\gamma_{t}s_{t}(\tilde{Z}_{t})dt+\sqrt{2\gamma_{t}}d\tilde{B}_{t}}_{\text{Langevin correction}};\;\tilde{Z}_{0}=Z_{0},$ (6)
for $t\in[0,1]$. This is obtained from adding a Langevin correction to the original RF ODE (3). Above, $\gamma_{t}\geq 0$ is a possibly time-varying diffusion coefficient, and $\{\tilde{B}_{t}\}_{t\geq 0}$ is the standard Brownian motion in $\mathbb{R}^{d}$. One can check (see, e.g., Liu (2024, Section 5.2) or Song et al. (2021b)) that the marginal $\tilde{p}_{t}:=\text{Law}(\tilde{Z}_{t})$ satisfies the continuity equation
$\frac{\partial\tilde{p}_{t}}{\partial t}+\nabla.(\tilde{v}_{t}\tilde{p}_{t})=0,$ (7)
where $\tilde{v}_{t}(x)=v_{t}(x)+\gamma_{t}s_{t}(x)-\gamma_{t}\nabla\log\tilde{p}_{t}(x)$. Since the continuity equations (4) and (7) coincide, $\tilde{p}_{t}=p_{t}$ , i.e., $\text{Law}(\tilde{Z}_{t})=\text{Law}(Z_{t})$, for all $t$, demonstrating that the SDE (6) can indeed be used for sampling from $p_{1}$. In practice, this is accomplished via an appropriate time discretization scheme
$\tilde{Z}_{t_{i+1}}=\tilde{Z}_{t_{i}}+\tilde{\eta}_{i}\left\{{v_{t_{i}}(\tilde{Z}_{t_{i}})+\gamma_{t_{i}}s_{t_{i}}(\tilde{Z}_{t_{i}})}\right\}+\sqrt{2\tilde{\eta}_{i}\gamma_{t_{i}}}\xi_{t_{i}},$ (8)
where $\left\{{t_{i}}\right\}_{i=0}^{N}$ is an appropriate schedule of times, the $\xi_{t_{i}}$’s are i.i.d. draws from $N(0,I_{d})$, and the $\tilde{\eta}_{i}$’s are (typically small) step-sizes. The Langevin term essentially acts as a course-corrector to the trajectory of (8) to adjust the trajectory distribution closer to $p_{t}$. Finally, one has freedom to choose the diffusion coefficients $\gamma_{t}$. In this paper we will restrict ourselves to the choice $\gamma_{t}=(1-t)/t$, which, as we will see below, is the appropriate noise scaling that yields equivalence between the STOC-RF SDE (6) and DDPM through the lens of Stochastic localization. This also addresses an question raised during the ICML tutorial talk Liu (2025) regarding the motivation of such choice.
3 Stochastic localization, RF and DDPM
Stochastic localization (SL).
SL is a probabilistic technique to study the properties of high-dimensional distributions due to Eldan (2013, 2020) that provides a powerful, unifying framework for understanding, analyzing, and improving score-based generative models and sampling algorithms. See El Alaoui and Montanari (2022), Montanari (2023) and the survey by Shi et al. (2025). To sample from the target data distribution $p_{1}$, SL focuses on an appropriate measured-valued stochastic process $\{\nu_{s}\}_{s\geq 0}$ such that $\nu_{s}$ “localizes” towards $p_{1}$ as $s\to\infty$. A straightforward way to formulate this process is by letting, for any $s\geq 0$, $\nu_{s}:=\text{Law}(X_{1}\mid U_{s})$, where $X_{1}\sim p_{1}$ and
$U_{s}=sX_{1}+B_{s},$ (9)
with $\{B_{s}\}_{s\geq 0}$ a standard Brownian motion in $\mathbb{R}^{d}$ (Montanari, 2023), independent of $X_{1}$. Thus, $\nu_{s}$ is a random measure depending on $U_{s}$. Note that $U_{s}/s\to X_{1}$ almost surely as $s\to\infty$, i.e., $\nu_{s}$ does indeed localize around $p_{1}$.
While enlightening, the construction (9) does not yield a sampler from $p_{1}$, as it uses $X_{1}\sim p_{1}$ in the definition of $U_{s}$. Therefore, it is instead convenient to consider the alternative process $\{\tilde{U}_{s}\}_{s\geq 0}$ defined as the solution of the SDE
$d\tilde{U}_{s}=a_{s}(\tilde{U}_{s})+d\tilde{B}_{s},$ (10)
[…]
Lemma 3.1.
For all $s\geq 0$, $\frac{d}{ds}\mathbb{E}\left[{\mathbf{A}_{s}}\right]=-\mathbb{E}\left[{\mathbf{A}_{s}^{2}}\right]$.
This is a crucial result for controlling the discretization error of ODE and SDE samplers of RF.
SL and DDPM.
DDPMs are a class of diffusion-based sampling mechanism for generative AI introduced by Ho et al. (2020). DDPM first starts with a forward process $(Y^{\prime}_{\tau})_{\tau\geq 0}$ starting from $Y_{0}^{\prime}=X_{1}\sim p_{1}$ and evolving according to the SDE
$dY_{\tau}^{\prime}=-\beta(\tau)Y_{\tau}^{\prime}d\tau+\sqrt{2\beta(\tau)}dB^{\prime}_{\tau},\;Y_{0}^{\prime}\sim p_{1},$ (11)
where $\beta(\tau)$ is the diffusion coefficient and $\left\{{B^{\prime}_{\tau}}\right\}_{\tau\geq 0}$ is standard Brownian motion. If $\beta(\tau)=1$, then the SDE (11) reduces to the classical Ornstein-Ulhenbeck (OU) process. Typically, The forwards process is continued till a large time $\tau=N$ such that $Y_{N}^{\prime}$ is approximately Gaussian. Let $q_{\tau}(y_{\tau}^{\prime})$ be the marginals of $y_{\tau}^{\prime}$ following (11). Then in order to sample from the target distribution $p_{1}$, a reverse process $\{\overleftarrow{Y}_{\tau}\}_{\tau\in[0,N]}$ is formulated, starting at $\overleftarrow{Y}_{0}=Y_{N}^{\prime}$ and evolving according to the reverse-time SDE
$d\overleftarrow{Y}_{\tau}=$ $\{\overleftarrow{Y}_{\tau}+2\nabla\log q_{N-\tau}(\overleftarrow{y}_{\tau})\}\beta(N-\tau)d\tau$ (12)
$+\sqrt{2\beta(N-\tau)}d\overleftarrow{B}_{\tau},$
where $\{\overleftarrow{B}_{\tau}\}_{\tau\geq 0}$ is a standard Brownian motion. Classical results in SDE (Anderson, 1982; Haussmann and Pardoux, 1986) show that $\overleftarrow{Y}_{\tau}\overset{d}{=}Y_{N-\tau}^{\prime}$ which allows us to generate samples from $p_{1}$ via simulating the backward SDE (12).
It is known that DDPMs and stochastic localization are equivalent (Montanari, 2023). In particular, Benton et al. (2023) showed that the forward process (11) is equivalent in law to the SL process (9) under an appropriate time change. In this section, we restate this result in a slightly different fashion that will allow us to easily see the connection between DDPM and RF. To this end, we define $\omega_{\tau}:=\exp(-2\int_{0}^{\tau}\beta(u)du)$. Then the following lemma shows that the the forward process (11) and the SL process (9) are equivalent:
Lemma 3.2.
Let $\{\tau(s)\}_{s\geq 0}$ be sequence such that $\frac{1-\omega_{\tau(s)}}{\omega_{\tau(s)}}=\frac{1}{s}$. Then, we have $\left\{{\frac{Y^{\prime}_{\tau(s)}}{\sqrt{\omega_{\tau(s)}}}}\right\}_{s\geq 0}\overset{d}{=}\left\{{\frac{U_{s}}{s}}\right\}_{s\geq 0}.$
If $\beta(\tau)=1$ for all $\tau$ (OU process), then the time transformation in Lemma 3.2 simplifies to be $\tau(s)=\frac{1}{2}\log(1+s^{-1})$, and we get $\left\{{\frac{Y^{\prime}_{\tau(s)}}{e^{-\tau(s)}}}\right\}_{s\geq 0}\overset{d}{=}\left\{{\frac{U_{s}}{s}}\right\}_{s\geq 0}$. Note, this matches the findings in Benton et al. (2023, Section 1.2). The proof is deferred to Appendix C.1.
SL, RF and STOC-RF.
Next, we will also show that the probability path of RF is also equivalent to the SL process (9). Recall that that RF starts by constructing the linear process $X_{t}=tX_{1}+(1-t)X_{0}$, where $X_{1}\sim p_{1},X_{0}\sim N(0,I_{d})$ and $t\in[0,1]$. To this end, we consider the process
$\tilde{X}_{t}=tX_{1}+tW_{(1-t)^{2}/t^{2}},\quad t\in[0,1]$ (13)
where $\{W_{\tau}\}_{\tau\geq 0}$ is a Brownian motion. Note that $\text{Law}(\tilde{X}_{t})=\text{Law}(X_{t})$, for all $t\in[0,1]$. In fact, as the next result show, the entire process (13) has the same law of the SL process (9), up to time-change.
Lemma 3.3.
Let $\{t(s)\}_{s\geq 0}$ be a sequence such that $\left({\frac{1-t(s)}{t(s)}}\right)^{2}=\frac{1}{s}$. Then, we have $\left\{{\frac{\tilde{X}_{t(s)}}{t(s)}}\right\}_{s\geq 0}\overset{d}{=}\left\{{\frac{U_{s}}{s}}\right\}_{s\geq 0}.$
The proof is in Appendix C.1. Crucially, Lemma 3.3 yields the following Corollary to Lemma 3.1 which is used to control the discretization error of the RF sampler.
Corollary 3.4.
Let ${\boldsymbol{\Sigma}}_{t}:=\text{Cov}(X_{1}\mid X_{t})$ for $t\in[0,1)$. Then $\frac{d}{dt}\mathbb{E}({\boldsymbol{\Sigma}}_{t})=-\frac{2t}{(1-t)^{3}}\mathbb{E}({\boldsymbol{\Sigma}}_{t}^{2})$.
See Section D.1 for complete proof. Additionally, Lemma 3.2 and Lemma 3.3 leads to the following:
$\left\{{\frac{Y^{{}^{\prime}}_{\tau}}{\sqrt{\omega_{\tau}}}}\right\}_{\tau\geq 0}\overset{d}{=}\left\{{\frac{\tilde{X}_{t(\tau)}}{t(\tau)}}\right\}_{\tau\geq 0},$ (14)
where $t(\tau)=\frac{\sqrt{\omega_{\tau}}}{\sqrt{\omega_{\tau}}+\sqrt{1-\omega_{\tau}}}$. In fact, under the same time-change the STOC-RF process (6) and the backward process (12) are equivalent.
Proposition 3.5.
Let $t(\tau)=\frac{\sqrt{\omega_{\tau}}}{\sqrt{\omega_{\tau}}+\sqrt{1-\omega_{\tau}}}$, and $\{\tilde{Z}_{t}\}_{t\geq 0}$ be a solution to the SDE (6). Then, $\left\{{\frac{\sqrt{\omega_{N-\tau}}\tilde{Z}_{t(N-\tau)}}{t(N-\tau)}}\right\}_{\tau\geq 0}$ is a solution to the SDE (12). Conversely, if $\overleftarrow{Y}_{\tau}$ is a solution of SDE (12), then $\left\{{\frac{\overleftarrow{Y}_{\tau(t)}}{\sqrt{\omega_{N-\tau(t)}}+\sqrt{1-\omega_{N-\tau(t)}}}}\right\}_{\tau\geq 0}$ is a solution of SDE (6), where $\tau(t)$ is the unique solution of the integral equation $\int_{0}^{N-\tau}\beta(u)du=\log\left({\sqrt{1+\frac{(1-t)^{2}}{t^{2}}}}\right)$.
The above proposition is very useful in analyzing the convergence rate of STOC-RF. Essentially, Proposition 3.5 allows uso to transit between DDPM and STOC-RF under simple time change and scaling. Therefore, convergence properties of one process can be carried over to the other one with the same transformation. This is the key insight that we will use to prove our convergence result of STOC-RF in later sections. The detailed proof is deferred to Appendix C.2.
SL and Stochastic Interpolants.
Lemma 3.3 and the equivalence result in (14) can be extended to cover general stochastic interpolation (Albergo and Vanden-Eijnden, 2023; Albergo et al., 2023) schemes beyond RF. In detail, instead of a linear interpolation, one can consider a general deterministic affine interpolation process $I_{\theta}=a_{\theta}X_{1}+b_{\theta}X_{0}$, where $a_{(\cdot)}$ and $b_{(\cdot)}$ are smooth functions of $\theta\in[0,1]$, with $a_{0}=b_{1}=0$, and $a_{1}=b_{0}=1$. When $X_{0}\sim N(0,I_{d})$, the associated stochastic interpolant is the process $(\tilde{I}_{\theta})_{\theta\in[0,1]}$, where $\tilde{I}_{\theta}=a_{\theta}X_{1}+a_{\theta}W_{r^{2}_{\theta}}$, with $r_{\theta}=b_{\theta}/a_{\theta}$ and $\left\{{W_{t}}\right\}_{t\geq 0}$ a Brownian motion. It is easy to see that $\tilde{I}_{\theta}\overset{d}{=}I_{\theta}$, for each $\theta$. Moreover, each stochastic interpolant can be rescaled and time-transformed to return the SL process, in a sense made precise by the next result.
Proposition 3.6.
$\left\{{\frac{\tilde{I}_{\theta(s)}}{a_{\theta(s)}}}\right\}_{s\geq 0}\overset{d}{=}\left\{{\frac{U_{s}}{s}}\right\}_{s\geq 0}$, with $r^{2}_{\theta(s)}=\frac{1}{s}$.
In particular, in virtue of Lemma 3.3, RF is equal in law to any interpolant process $I_{\theta}$ under the time change $t(\theta)=\frac{a_{\theta}}{a_{\theta}+b_{\theta}}$. This equivalence among stochastic interpolants is not new, as it has also been shown in Liu (2024) by direct arguments. Connection between diffusion and flow matching has also been studied in Albergo et al. (2023); Albergo and Vanden-Eijnden (2022); Gao et al. (2024); Ma et al. (2024); Rout et al. (2025) through different perspectives. By leveraging the representation properties of the SL process, we establish this result in a more unified manner, which has the added benefit of directly applying to both stochastic interpolants and SDEs. The proof of Proposition 3.6 is deferred to Appendix C.3.
4 Main Results
In this section, we present our main convergence results for both deterministic and stochastic samplers. We begin by introducing and commenting on our key assumptions. Throughout, $\left\{{t_{i}}\right\}_{i=0}^{N}$ refers to the time schedule for both the RF and STOC-RF procedures, as in (5) and (8), respectively, and $N$ is the number of time steps.
4.1 Low-dimensionality of target distribution
To formalize the notion of low-dimensionality of $p_{1}$, we turn to the geometric complexity of its support $\mathcal{X}$ as measured by its metric entropy.
Assumption 4.1 (Low-dimensionality).
Let $\epsilon_{0}=N^{-c_{\epsilon_{0}}}$ for some sufficiently large $c_{\epsilon_{0}}>0$. The $\epsilon_{0}$ covering number of $\mathcal{X}$ with respect to the Euclidean norm, $\mathcal{N}(\mathcal{X},\|\cdot\|_{2},\epsilon_{0})$, satisfies $\log\mathcal{N}(\mathcal{X},\|\cdot\|_{2},\epsilon_{0})\leq C_{\rm cover}k\log N,$ for some constant $C_{\rm cover}>0$ and integer $k$. The quantity $k$ is referred as the intrinsic-dimension of $\mathcal{X}$.
The above assumption was introduced in recent works by Huang et al. (2024); Li and Yan (2024); Liang et al. (2025) on adaptation of diffusion methods to intrinsic low dimensionality of the target function. Notably, by increasing the number of steps $N$, the assumption is less and less stringer. Informally, Assumption 4.1 requires the support of $p_{1}$ to be concentrated on or near a $k$-dimensional set. The above definition of intrinsic dimension is fairly general, as it covers a variety of important low-dimensional structures. Examples satisfying Assumption 4.1 include $k$-dimensional linear subspace in $\mathbb{R}^{d}$ and $k$-dimensional non-linear manifolds (provided $\mathcal{X}$ is polynomially bounded as in Assumption 4.2). Structures with doubling-dimension $k$ (Dasgupta and Freund, 2008; Kpotufe and Dasgupta, 2012) also satisfies Assumption 4.1. More detailed discussion on this can be found in Section 4.1.1 of Huang et al. (2024). Moreover, in practice, target distribution (e.g., image data) is supported on a lower dimensional manifold and the intrinsic dimension $k\ll d$ (Facco et al., 2017; Ansuini et al., 2019; Pope et al., 2021).
Assumption 4.2 (Bounded support).
There exists a universal constant $c_{R}>0$ such that $\sup_{x\in\mathcal{X}}\|x\|_{2}\leq R,\text{where $R=N^{c_{R}}$}.$
The bounded support assumption is not as stringent as it may appear at first, as the diameter of $\mathcal{X}$ is allowed to scale polynomially (with arbitrarily large degree) in the number of iterations of the sampler, which accommodates a wide range of practical applications, such as image generation.
4.2 Assumptions on velocity approximation error
Our next set of assumptions are concerned with the approximation error of the estimated velocity field $\hat{v}_{t}$.
Assumption 4.3 (Velocity approximation).
Define $(\varepsilon^{\rm v}_{i})^{2}:=\mathbb{E}\|v_{t_{i}}(X_{t_{i}})-\widehat{v}_{t_{i}}(X_{t_{i}})\|_{2}^{2}$. Then, $\frac{1}{N}\sum_{i=1}^{N}(\varepsilon^{\rm v}_{i})^{2}\leq\varepsilon_{\rm v}^{2}.$
The above assumption essentially requires a good control on the mean-squared error of the estimated velocity field. Similar assumptions have been considered in Bansal et al. (2025); Guan et al. (2025) to prove Wasserstein convergence of RF. In the context of score-matching algorithms, similar assumptions related to the score-approximation error have also appeared in prior works such as Benton et al. (2023); Chen et al. (2023); Li et al. (2024b); Liang et al. (2025).
Assumption 4.4 (Higher-order approximations).
  1. (a)
    Let $(\varepsilon^{\rm J,1}_{i})^{2}:=\mathbb{E}\|\nabla v_{t_{i}}(X_{t_{i}})-\nabla\widehat{v}_{t_{i}}(X_{t_{i}})\|_{F}^{2}$. Then, $N^{-1}\sum_{i=1}^{N}(\varepsilon^{\rm J,1}_{i})^{2}\leq\varepsilon_{\rm J,1}^{2}$. Also, we assume $\sup_{x\in\mathbb{R}^{d}}\eta_{i}\|\nabla v_{t_{i}}(x)-\nabla\widehat{v}_{t_{i}}(x)\|_{{\rm op}}\leq 1/8$ for all $0\leq i\leq N-2$.
  2. (b)
    Let $(\varepsilon^{\rm J,2}_{i})^{2}:=\mathbb{E}[\operatorname{Tr}(\nabla v_{t_{i}}(X_{t_{i}})-\nabla\widehat{v}_{t_{i}}(X_{t_{i}}))^{2}]$. Then, $N^{-1}\sum_{i=1}^{N}(\varepsilon^{\rm J,2}_{i})^{2}\leq\varepsilon_{\rm J,2}^{2}$.
  3. (c)
    Let $(\varepsilon^{\rm H}_{i})^{2}:=\mathbb{E}\|\nabla\operatorname{Tr}(\nabla v_{t_{i}}(X_{t_{i}})-\nabla\widehat{v}_{t_{i}}(X_{t_{i}}))\|_{2}^{2}$. Then, $N^{-1}\sum_{i=1}^{N}(\varepsilon^{\rm J,2}_{i})^{2}\leq\varepsilon_{\rm H}^{2}$.
Nearly identical assumptions have been used in Liang et al. (2025) in the context of DDIM. Intuitively, an adequate control of the higher-order approximation terms mitigates the propagation of error during sampling, as the deterministic samplers are unable to self-correct the path of the flow.
4.3 Convergence rate of rectified flow
Time partition.
Our convergence guarantees for the RF sampler (5) depend crucially on a carefully constructed U-shaped time schedule $\{t_{i}\}_{i=0}^{N}$, which we describe next. For given, user-specified values $h\in(0,1)$ and $\delta\geq 0$, we consider the non-uniform time partition
$t_{j}=\begin{cases}0,&j=0,\\[4.0pt] \delta,&j=1,\\[6.0pt] (1+h)\,t_{j-1},&2\leq j\leq\dfrac{N}{2},\\[8.0pt] 1-(1+h)\,(1-t_{j+1}),&\dfrac{N}{2}<j\leq N-2,\\[6.0pt] 1-\delta,&j=N-1,\\[6.0pt] 1,&j=N.\end{cases}$ (15)
We set $h$ such that $t_{N/2}=\delta(1+h)^{(N-2)/2}=1/2$. Using N2>N/2N-2>N/2 (for $N>4$), we arrive at
$\frac{h}{2}\leq\log(1+h)=\frac{2\log(\frac{1}{2\delta})}{N-2}\Rightarrow h\leq\frac{8\log(\frac{1}{2\delta})}{N}.$ (16)
We remark that the proposed time schedule is practicable, as it does not depend on any properties of $p_{1}$. Typically, $\delta$ is chosen to be small, as this give rise to a U-shaped time discretization as shown in Figure 2. This type of choice is not unusual (Lee et al., 2024), as it yields more accurate estimates $\widehat{v}_{t}$ during training when $t\approx 0$ and $t\approx 1$. Prior empirical work by non-uniform time discretizations have also been considered by Huang et al. (2024); Li and Yan (2024); Liang et al. (2025) in the context DDPM and DDIM to establish low-dimensional adaptation of the models. It is important to emphasize that the specific form of the time schedule (15), while conforming to best empirical practices, is theoretically grounded, and it was derived directly as a result of our technical analysis of the RF dynamics (see Step 7 of the proof of Theorem 4.5, Appendix B.2). We now present the main convergence result for the RF sampler.
Figure 2: Histogram and kernel density estimation (KDE) plot of time-grid (15) showing U-shaped distribution.
Theorem 4.5.
Let Assumptions 4.1-4.4 hold, and write $\mu_{1}=\mathbb{E}_{X\sim p_{1}}[X]$. Let $\{Y_{t_{i}}\}_{i\geq 0}$ denote the RF updates of (5) with time schedule (15). Then, for $\delta=1/(N\vee d)$, we have
${\sf TV}(p_{X_{t_{N-1}}},p_{Y_{t_{N-1}}})$ (17)
$\lesssim\frac{k\log^{3}(\frac{1}{\delta})}{N}+\frac{\log^{2}(\frac{1}{\delta})}{N}\varepsilon_{\rm J,1}^{2}+\frac{\log^{2}(\frac{1}{\delta})}{N}\varepsilon_{\rm v}^{2}$
$\quad+\log\left({1/\delta}\right)(\varepsilon_{\rm v}+\varepsilon_{\rm J,1}+\varepsilon_{\rm J,2}+\varepsilon_{\rm H})$
$\quad+\delta^{2}\mathbb{E}\|X_{1}-\mu_{1}\|_{2}^{2}+\delta^{2}(\varepsilon^{\rm v}_{0})^{2}$
$\quad+\delta^{2}(\varepsilon^{\rm J,1}_{0})^{2}+(\varepsilon^{\rm v}_{0}+\varepsilon^{\rm J,1}_{0}+\varepsilon^{\rm J,2}_{0}+\varepsilon^{\rm H}_{0})+\delta^{-2}N^{-10}.$
Adaptation to low-dimensionality and accelerated convergence. The rate in Theorem 4.5 does not contain a direct dependence on the ambient dimension $d$ and instead scales linearly in the intrinsic dimension $k$ of $\mathcal{X}$, indicating that the RF sampler adapts to the low-dimensional structure of the target distribution. Thus, in principle, assuming perfect knowledge of the velocity drift (i.e. $\widehat{v}_{t}=v_{t}$ and the higher approximation terms of 4.4 are all zero), the RF sampler (5) needs $\tilde{O}(k/\epsilon)$ iterations to achieve the guarantee ${\sf TV}(p_{X_{t_{N-1}}},p_{Y_{t_{N-1}}})<\epsilon$. If the intrinsic dimension $k\ll d$, then the RF sampler automatically accelerates without any prior knowledge of the low-dimensionality of $p_{1}$. To the best of our knowledge, this is the first work that provides a convergence guarantee of RF that adapts to the low-dimensionality of $p_{1}$.
The dependence on $\mathbb{E}\|X_{1}-\mu_{1}\|_{2}^{2}$. The upper bound in Theorem 4.5 contains a term involving $\mathbb{E}\|X_{1}-\mu_{1}\|_{2}^{2}$, which even for data supported on manifolds could scale linearly in the ambient dimension $d$. However, in practice one can choose the free parameter $\delta$ to be e.g. $O(d^{-1})$, so that the term becomes negligible. With this choice of $\delta$, last term in (17) scales as $d^{2}N^{-10}$ which is also negligible in practice.
Guarantee on perturbed data. The TV guarantee in Theorem 4.5 is on the penultimate update $Y_{t_{N-1}}$ of (5), and not on $Y_{t_{N}}=Y_{1}$. It is worthwhile to mention that ${\sf TV}(p_{X_{1}},p_{Y_{t_{N}}})$ might not be a useful or even meaningful quantity, as $X_{1}$ and $Y_{1}$ may have supports of different dimensions. For example, $X_{1}$ could be supported on a low-dimensional space, whereas $Y_{t_{N}}$ would typically have full dimensional support, thereby rendering ${\sf TV}(X_{1},Y_{t_{N}})=1$. However, note that for any $\delta>0$, $X_{\delta}\overset{d}{=}(1-\delta)X_{1}+\delta X_{0}$ has full-dimensional support due to the slight Gaussian perturbation. Therefore, the TV distance in Theorem 4.5 is well defined.
Higher-order approximation error terms. The upper bound in Theorem 4.5 is heavily dependent on the higher-order approximation errors of $\widehat{v}_{t}$. These terms are required to be small as deterministic samplers are unable to self-correct their trajectories. In addition, at $t=0$, the true drift $v_{0}(x)=\mu_{1}-x$ can be estimated by $\widehat{v}_{0}(x)=n^{-1}\sum_{i\in[n]}{X_{1}^{(i)}}-x$, where $\{X_{1}^{(i)}\}_{i\in[n]}\overset{i.i.d.}{\sim}p_{1}$. With this choice of $\widehat{v}_{0}$, we have $\varepsilon^{\rm J,1}_{0}=\varepsilon^{\rm J,2}_{0}=\varepsilon^{\rm H}_{0}=0$, and $\varepsilon^{\rm v}_{0}=\sqrt{\mathbb{E}\|X_{1}-\mu_{1}\|_{2}^{2}/n}$.
4.4 Convergence rate of STOC-RF
In this section, we will leverage the equivalence between DDPM and STOC-RF to design an SDE-based sampler that also adapts to the low-dimensionality of $p_{1}$. We begin with the solution $\left\{{Y^{\prime}}\right\}_{\tau\geq 0}$ to the forward process (11). We also consider the time-change map $\tau\in[0,\infty)\mapsto t(\tau)=\frac{\sqrt{\omega_{\tau}}}{\sqrt{\omega_{\tau}}+\sqrt{1-\omega_{\tau}}}.$ In light of (14) and Proposition 3.5, we note that $\left\{{Y^{\prime}_{\tau}}\right\}_{\tau\geq 0}$ and $\{\tilde{Z}_{t}\}_{t\in[0,1]}$ (recall 6) are equivalent under the time change $\tau\mapsto t(\tau)$, and we have $\frac{Y^{\prime}_{\tau}}{\sqrt{\omega_{\tau}}}\overset{d}{=}\frac{\tilde{Z}_{t(\tau)}}{t(\tau)}$. In practice, we simulate the forward process by discretizing $\tau$ to positive integers and setting specific values for $\omega_{\tau}$. Here, we let $\omega_{\tau}=\prod_{j=1}^{\tau}\alpha_{j}$ with $\alpha_{j}$’s defined implicitly as
$\beta_{1}=1-\alpha_{1}=1/N^{c_{0}},$ (18)
$\beta_{\tau+1}=1-\alpha_{\tau+1}$
$=\frac{c_{1}\log N}{N}\min\left\{{\beta_{1}\left({1+\frac{c_{1}\log N}{N}}\right)^{\tau},1}\right\},\quad 1\leq\tau\leq N,$
for a user-specified choice of the parameters $c_{0},c_{1}>0$. The time schedule above is very similar to the ones considered by Potaptchik et al. (2024); Li et al. (2024b); Li and Yan (2024); Huang et al. (2024); Liang et al. (2025) to analyze the convergence properties of DDPM. Next, we recall the DDPM sampler considered in Liang et al. (2025); Huang et al. (2024); Li and Yan (2024):
$\hat{Y}^{\prime}_{\tau-1}=\frac{1}{\sqrt{\alpha_{\tau}}}\left\{{\hat{Y}^{\prime}_{\tau}+\delta_{\tau}\widehat{s}_{Y^{\prime}_{\tau}}(\hat{Y}^{\prime}_{\tau})+\nu_{\tau}\xi_{\tau}}\right\},\hat{Y}^{\prime}_{N}\sim N(0,I_{d})$ (19)
$\delta_{\tau}=1-\alpha_{\tau},\nu_{\tau}=\sqrt{\frac{(\alpha_{\tau}-\omega_{\tau})(1-\alpha_{\tau})}{1-\omega_{\tau}}},$
$\tau=N,N-1,\ldots,1,\quad\{\xi_{\tau}\}_{\tau\geq 1}\overset{i.i.d.}{\sim}N(0,I_{d}),$
where $\widehat{s}_{Y^{\prime}_{\tau}}$ is an approximate score of $Y^{\prime}_{\tau}$. To provide some intuition behind this sampler, we recall that $1-\alpha_{\tau}=\beta_{\tau}\approx 0$. Hence, $1/\sqrt{\alpha_{\tau}}\approx 1+\beta_{\tau}/2,\delta_{\tau}/\sqrt{\alpha_{\tau}}\approx\beta_{\tau}/2$, and $\nu_{\tau}/\sqrt{\alpha_{\tau}}\approx\sqrt{\beta_{\tau}}$. Substituting this in (19), we get
$\hat{Y}^{\prime}_{\tau-1}\approx\hat{Y}^{\prime}_{\tau}+\frac{\beta_{\tau}}{2}\left\{{\hat{Y}^{\prime}_{\tau}+2\widehat{s}_{Y^{\prime}_{\tau}}(\hat{Y}^{\prime}_{\tau})}\right\}+\sqrt{2\left({\frac{\beta_{\tau}}{2}}\right)}\xi_{\tau},$ (20)
which is essentially the discretized version of (12) with step-size $\beta_{\tau}/2$. Therefore, the sampler (19) roughly approximates the DDPM backward process.
Next, we convert (19) to a STOC-RF sampler using the equivalence between $Y^{\prime}_{\tau}$ and $\tilde{Z}_{t(\tau)}$. To this end, we let, for any $t\in[0,1]$, $\sigma_{t}^{2}=(1-t)^{2}+t^{2}$, so that
$\sigma_{t(\tau)}^{2}=\frac{1}{(\sqrt{\omega(\tau)}+\sqrt{1-\omega(\tau)})^{2}}=\frac{t^{2}(\tau)}{\omega_{\tau}}.$
Therefore, we have $Y^{\prime}_{\tau}\overset{d}{=}\tilde{Z}_{t(\tau)}/\sigma_{t(\tau)}$, and $s_{Y^{\prime}_{\tau}}(y)=\sigma_{t(\tau)}s_{t(\tau)}(\sigma_{t(\tau)}y)$. We also define $t_{i}:=t(N-i)$ for $i=0,1,\ldots,N-1$, and $R_{i}=t_{i}/\sigma_{t_{i}}$. Collecting everything above, we arrive at an equivalent reformulation of (19) for RF sampling of the form
$\frac{Y_{t_{i+1}}}{\sigma_{t_{i+1}}}=\frac{R_{i+1}}{R_{i}}\left\{{\frac{Y_{t_{i}}}{\sigma_{t_{i}}}+\eta_{i}\sigma_{t_{i}}\widehat{s}_{t_{i}}(Y_{t_{i}})+\sqrt{\psi_{i}}W_{t_{i}}}\right\},$ (21)
$\frac{Y_{t_{0}}}{\sigma_{t_{0}}}\sim N(0,I_{d}),\quad\{W_{t_{i}}\}_{i\geq 0}\overset{i.i.d.}{\sim}N(0,I_{d}),$
$i=0,1,\ldots,N-1,$
where $\eta_{i}=1-\frac{R_{i}^{2}}{R_{i+1}^{2}},\psi_{i}=\frac{R_{i}^{2}}{R_{i+1}^{2}}\frac{1-R_{i+1}^{2}}{1-R_{i}^{2}}\left({1-\frac{R_{i}^{2}}{R_{i+1}^{2}}}\right)$, and $\widehat{s}_{t_{i}}(x)=\frac{t_{i}\widehat{v}_{t_{i}}(x)-x}{1-t_{i}}$. See Appendix B.3 for details. Lastly, we impose an assumption on the approximation error of $\widehat{s}_{t_{i}}$.
Assumption 4.6.
Define $(\varepsilon^{\rm s}_{i})^{2}:=\mathbb{E}\|s_{t_{i}}(X_{t_{i}})-\widehat{s}_{t_{i}}(X_{t_{i}})\|_{2}^{2}$. Then, $\frac{1}{N}\sum_{i=1}^{N}(\varepsilon^{\rm s}_{i})^{2}\leq\varepsilon_{\rm s}^{2}.$
Leveraging the equivalence between (19) and (21), we obtain convergence guarantee for STOC-RF, which the second main result of this paper.
Theorem 4.7.
Under Assumption 4.1, 4.2 and 4.6, the stoc-RF sampler (21) satisfies
${\sf TV}(Y_{t_{N-1}},X_{t_{N-1}})\lesssim\frac{k\log^{3}N}{N}+\varepsilon_{\rm s}\sqrt{\log N}.$
Note that $t_{N-1}=\frac{\sqrt{\omega_{1}}}{\sqrt{\omega_{1}}+\sqrt{1-\omega_{1}}}\approx 1$. Therefore, $\text{Law}(X_{t_{N-1}})\approx p_{1}$. In comparison with Theorem 4.5, STOC-RF sampler does not require any conditions on higher-order derivatives (4.4) of the score or velocity field. That is, STOC-RF sampler can achieve similar performance under less stringent requirements. This is primarily due to the Gaussian perturbation term in (21) which acts as a corrector step. The proof is in Appendix B.3.
Remark 4.8.
Using the equivalence relation (14), it is also possible to directly relate the deterministic RF sampler to the DDIM sampler studied by Liang et al. (2025, Theorem 1) (Appendix B.4) via the same time change transformation considered in this section. The resulting time schedule also follows a U-shaped distribution and the sampler will also adapts to the low-dimensionality of $p_{1}$.
5 Experiments
We perform extensive simulations to corroborate our theoretical findings for both synthetic data and T2I experiments. Codes are available in https://github.com/roysaptaumich/Low-dim-RF.
Figure 3: (a)-(b) Estimated TV  distance between blurred low-rank Gaussian distribution and the generated samples by the RF sampler (5) under uniform and non-uniform time-grid under varying $d$. (c)-(d) Estimated TV  distance between blurred low-rank Gaussian distribution and the generated samples by the RF and STOC-RF samplers under uniform and non-uniform time-grid under varying $d$.
(a) Prompt: Grape vines in the shape of text ‘open your mind’ sprouting out of a head with flowers and butterflies. DSLR photo. (Better text rendering, $N=50$)
(b) Prompt: A sign that says ”Please refrain from arguing with the chimpanzees”. (Minimal mistakes, $N=50$)
(c) Prompt: Photo of a robot lecturer writing the words ‘Representation Learning’ in cursive on a blackboard, with math formulas and diagrams. (Correct spelling, $N=50$)
Figure 4: Generated images using Flux for specific prompts via Euler’s scheme (5) under non-uniform time-grid (15) (Red frame) and uniform time-grid (Blue frame). Generation quality of RF is better under non-uniform time-grid (15): less hallucinations, enjoys better text rendering and achieves better semantic completeness.
Synthetic data. In this section we demonstrate the usefulness of non-uniform time-grid for generation task from low-dimensional distribution. We set the intrinsic dimension $k=8$, and set our target distribution to be $p_{1}=N(\mu,\Sigma)$, where $\mu=(8,8,\ldots,8)^{\top}\in\mathbb{R}^{d}$ and $\Sigma$ is a block-diagonal matrix with blocks $I_{k}$ and $\boldsymbol{0}_{(d-k)\times(d-k)}$. Tweedie’s formula (Section 2.1) yields
$v_{t_{i}}(x)=\begin{bmatrix}\frac{2t_{i}-1}{\sigma_{t_{i}}^{2}}I_{k}&\boldsymbol{0}_{k\times(d-k)}\\ \boldsymbol{0}_{(d-k)\times k}&-\frac{1}{1-t_{i}}I_{d-k}\end{bmatrix}x.$ (22)
We use (22) for generating new samples from $p_{1}$. We vary $d$ between 10 and 800, and set $N\in\left\{{100,200}\right\}$. We use both a non-uniform time-grid (15) (with $\delta=\min\{1/N,1/d\}$) and a uniform time-grid for the generation task. We generate 2000 samples using (15) and measure the TV distance with respect to another batch of 2000 samples independently generated from the slightly blurred target distribution $\text{Law}((1-\delta)X_{1}+\delta Z)$, where $Z\sim N(0,I_{d})$. We use the discriminative approach proposed in Tao et al. (2024) to estimate the TV distance between the distributions via 10 independent Monte Carlo rounds. Figure 3(a)-(b) show that nonuniform discretization has better adaptation to the low-dimensional structure of $p_{1}$ for large $d$. For the SDE sampler (21), we also compare its TV convergence with that of the deterministic sampler (5) with the non-uniform time grid (15). Figure 3(c)-(d) show that both samplers have similar performance across varying dimensions.
T2I experiments. We perform experiments using Flux (Labs, 2024; Labs et al., 2025) to demonstrate the usefulness of the U-shaped time schedule (15). Figure 4 shows that generation quality of RF is superior under a non-uniform time-grid (15) compared to uniform time-grid. Figure 5 compares the generation quality of STOC-RF sampler (21) and the vanilla SDE sampler (20) and ODE samplers. We see that the generation quality is generally better for STOC-RF. Details are deferred to Appendix A.
(a) Prompt: A photo of a cat holding a sign Rectified flow. ($N=50$)
(b) Prompt: Grape vines in the shape of text ‘open your mind’ sprouting out of a head with flowers and butterflies. DSLR photo. ($N=50$)
(c) Prompt: Photo of a robot lecturer writing the words ‘Representation Learning’ in cursive on a blackboard, with math formulas and diagrams. ($N=50$)
(d) Prompt: A car made of cheese on a wooden table. ($N=80$)
Figure 5: Generated images using Flux for specific prompts via Euler’s scheme (5) under non-uniform time-grid (15) (Columns 1), uniform time-grid (Column 2), STOC-RF sampler (Column 3), and vanilla SDE sampler (Column 4).
Caution: The TV convergence rates in both Theorem 4.5 and Theorem 4.7 depend highly on the approximation errors. Difficulty may arise in estimating $v_{t}(\cdot)$ along the direction of low-dimensionality when $t\approx 1$ due to non-smoothness (see (22)). Such non-smoothness is usually hard to capture via standard neural networks (e.g., MLP, U-Net) that are typically smooth, and necessitate the adoption of more expressive architecture. We defer this to future research.
6 Conclusions
In this paper, we design a new time-discretization scheme that allows any off-the-shelf pretrained RF sampler to achieve accelerated convergence by adapting to the low-dimensional structure of the target distribution. This allows the RF model to achieve better sampling quality in the first reflow step itself and also corroborates the empirical findings in Lee et al. (2024). Next, we show that the DDPM is equivalent to a stochastic version of RF by establishing a novel connection between these processes and the stochastic localization. Building on this connection, we further design a stochastic RF sampler that also adapts to the low-dimensionality of the target distribution under less stringent assumptions on the learned RF model.
Acknowledgements
We thank Prof. Eric Vanden-Eijnden at NYU, Prof. Sanjay Shakkottai at UT Austin, and Dr. Christian Rau for helpful discussions and pointing out important references. AR, PS and SR gratefully acknowledge the NSF grants CCF-2019844 and CCF-2505865. PS and SR gratefully acknowledges NSF grant 2217069.
Appendix A Additional Experiments
We have performed various experiments to demonstrate the usefulness of the U-shaped time schedule (15) in the context of T2I. We use the Flux model (Labs, 2024; Labs et al., 2025) which is a pretrained RF model. For our experiments, we use the prompts designed by Liu et al. (2023); Hu et al. (2025b) to test and assess the generation quality. In these examples, we always set $\delta=1/N$ for a specified choices of $N$. In Figure 6 we see that generation quality of RF is superior under a non-uniform time-grid (15) compared to uniform time-grid. In particular, the experiments indicate that a non-uniform grid helps to reduce hallucinations, enjoys better text rendering and achieves better semantic completeness.
We also compare the generation quality of STOC-RF sampler (21) and the vanilla SDE sampler (20) in Figure LABEL:fig:_Flux_SDE/RF_experiments. The generated plots show that the quality is generally better for STOC-RF sampler (21) compared to the RF sampler and the vanilla SDE sampler (20).
(a) Prompt: A photo of a cat holding a sign Rectified flow. (Less hallucination, $N=50$)
(b) Prompt: Grape vines in the shape of text ‘open your mind’ sprouting out of a head with flowers and butterflies. DSLR photo. (Better text rendering, $N=50$)
(c) Prompt: the hubble telescope and the milky way, with the text “the universe is a mystery, but we are here to solve it” (Better fidelity/ prompt adherence, $N=50$)
(d) Prompt: A sign that says ”Please refrain from arguing with the chimpanzees”. (Minimal mistakes, $N=50$)
(e) Prompt: Photo of a robot lecturer writing the words ‘Representation Learning’ in cursive on a blackboard, with math formulas and diagrams. (Correct spelling, $N=50$)
(f) Prompt: A car made of cheese on a wooden table. (Semantic completeness, $N=80$)
Figure 6: Generated images using Flux for specific prompts via Euler’s scheme (5) under non-uniform time-grid (15) (Red frame) and uniform time-grid (Blue frame).
Appendix B Proof of main results
B.1 Velocity function and its first order identities
Before proceeding, we illustrate the connection between the velocity $v_{t}(x)$ and the score function $s_{t}(x):=\nabla\log p_{t}(x)$. To this end, we define two key quantities:
$\mu_{1\mid t}(x)=\mathbb{E}[X_{1}\mid X_{t}=x],\;\text{Cov}_{1\mid t}(x)=\text{Cov}(X_{1}\mid X_{t}=x).$
With this, true drift be re-written as
$v_{t}(x)=\mathbb{E}[X_{1}-X_{0}\mid X_{t}=x]=\frac{1}{1-t}\left({\mu_{1\mid t}(x)-x}\right).$ (23)
By the first and second order Tweedie’s formula (Efron, 2011; Meng et al., 2021), it follows that, for any choice of $t\in(0,1]$ and $x\in\mathbb{R}^{d}$,
$\mu_{1\mid t}(x)=\frac{x}{t}+\frac{(1-t)^{2}}{t}s_{t}(x),$
$\text{Cov}_{1\mid t}(x)=\frac{(1-t)^{2}}{t^{2}}I_{d}+\frac{(1-t)^{4}}{t^{2}}\nabla s_{t}(x).$
Substituting this in (23), we obtain that
$v_{t}(x)=\frac{x}{t}+\frac{(1-t)}{t}s_{t}(x),\quad\text{and}\quad\nabla v_{t}(x)=-\frac{I_{d}}{1-t}+\frac{t}{(1-t)^{3}}\text{Cov}_{1\mid t}(x).$
B.2 Proof of convergence of RF sampler (Theorem 4.5)
Notational conventions.
We begin with some notational conventions for the ease of presentation.
  1. 1.
    We define transition map for the Euler’s scheme as
    $\Phi_{t_{i}}(x):=x+\eta_{i}\widehat{v}_{t_{i}}(x)\quad\text{with $\eta_{i}=t_{i+1}-t_{i}$},$
    where $\widehat{v}_{t}(x)$ is the velocity estimate.
  2. 2.
    For notational brevity, we will use index $i$ to denote the time $t_{i}$. In this index notation we have $X_{N}\equiv X_{1}\sim p_{1}$
  3. 3.
    With the above notations in mind, we have
    $X_{i}=(1-t_{i})X_{0}+t_{i}X_{N},\quad\text{and}\quad Y_{i+1}=\Phi_{i}\left(Y_{i}\right).$ (24)
Step 1: TV recursion.
We start by obtaining a recursion relation between the TVdistances between the updates $Y_{i}$ and the ground truth $X_{i}$. For any $i\leq N-2$, we have
${\sf TV}\left(p_{X_{i+1}},p_{Y_{i+1}}\right)$ $=\sup_{\mathcal{A}\subseteq\mathbb{R}^{d}}\left\{\mathbb{P}_{X_{i+1}}(\mathcal{A})-\mathbb{P}_{Y_{i+1}}(\mathcal{A})\right\}=\sup_{\mathcal{A}\subseteq\mathbb{R}^{d}}\left\{\mathbb{P}_{X_{i+1}}(\mathcal{A})-\mathbb{P}_{Y_{i}}\left(\Phi_{i}^{-1}(\mathcal{A})\right)\right\}$ (25)
$\leq\sup_{\mathcal{A}\subseteq\mathbb{R}^{d}}\left\{\mathbb{P}_{X_{i+1}}(\mathcal{A})-\mathbb{P}_{X_{i}}\left(\Phi_{i}^{-1}(\mathcal{A})\right)\right\}+\sup_{\mathcal{A}\subseteq\mathbb{R}^{d}}\left\{\mathbb{P}_{X_{i}}\left(\Phi_{i}^{-1}(\mathcal{A})\right)-\mathbb{P}_{Y_{i}}\left(\Phi_{i}^{-1}(\mathcal{A})\right)\right\}$
$\leq\sup_{\mathcal{A}\subseteq\mathbb{R}^{d}}\left\{\mathbb{P}_{X_{i+1}}(\mathcal{A})-\mathbb{P}_{\Phi_{i}\left(X_{i}\right)}(\mathcal{A})\right\}+{\sf TV}\left(p_{X_{i}},p_{Y_{i}}\right)$
$={\sf TV}\left(p_{X_{i+1}},p_{\Phi_{i}\left(X_{i}\right)}\right)+{\sf TV}\left(p_{X_{i}},p_{Y_{i}}\right)$
where the first identity arises from the basic property of the TV distance.
Step 2: Simplifying ${\sf TV}(p_{X_{i+1}},p_{\Phi_{i}(X_{i})})$.
Next, we recall that $v_{i}(x)=\frac{x}{t_{i}}+\frac{1-t_{i}}{t_{i}}s_{i}(x)$. Therefore, by Liu (2024, Lemma 13) we have
$\frac{\partial v_{i}(x)}{\partial x}=-\frac{1}{1-t_{i}}I+\frac{t_{i}}{(1-t_{i})^{3}}\text{Cov}_{1\mid t_{i}}(x_{i})$ (26)
Further, for any vector $u\in\mathbb{R}^{d}$ with $\|u\|_{2}=1$ Assumption 4.4(a) tells us that for $i\leq N-2$
$u^{\top}\frac{\partial\Phi_{i}(x)}{\partial x}u=u^{\top}\left(I+\eta_{i}\frac{\partial\widehat{v}_{i}(x)}{\partial x}\right)u=u^{\top}\left(I+\eta_{i}\frac{\partial v_{i}(x)}{\partial x}\right)u+\eta_{i}u^{\top}\varepsilon_{i}^{J}(x)u$
$\geq u^{\top}\left(I+\eta_{i}\frac{\partial v_{i}(x)}{\partial x}\right)u-\frac{1}{8}\|u\|_{2}^{2}=u^{\top}\left\{\left(1-\frac{\eta_{i}}{1-t_{i}}\right)I+\frac{\eta_{i}t_{i}}{(1-t_{i})^{3}}\operatorname{Cov}_{1\mid t_{i}}(x)\right\}u-\frac{1}{8}\|u\|_{2}^{2}$
$\geq\left(1-h-\frac{1}{8}\right)\|u\|_{2}^{2}>\frac{3}{8},\quad\text{if $h<1/2$.}$
Here, the penultimate relation follows from the positive semi-definiteness of the covariance matrix $\operatorname{Cov}_{1\mid t}(x)$ and the fact that $\frac{\eta_{i}}{1-t_{i}}=\frac{ht_{i}}{1-t_{i}}{\mathbbm{1}}\left\{{t_{i}\leq 1/2}\right\}+\frac{h(1-t_{i+1})}{1-t_{i}}{\mathbbm{1}}\left\{{t_{i}>1/2}\right\}\leq h$ (see Lemma D.2). Here $\mathbbm{1}$$\left\{{\cdot}\right\}$ denotes the indicator function.
Invertibility of $\Phi_{i}(\cdot)$: The above inequality implies that $\frac{\partial{\Phi}_{i}(x)}{\partial x}$ is positive definite uniformly over all $x$. This entails that $\Phi_{i}(\cdot)$ is proper, i.e, $\lim_{\|x\|_{2}\to\infty}\|\Phi(x)\|_{2}=\infty$. To see this, we define $\phi(t):=u^{\top}\Phi_{i}(tu)$ for a fixed unit vector $u$.Then, we have $\phi^{\prime}(t)=u^{\top}\nabla\Phi_{i}(tu)u>3/8$. This shows that
$\phi(t)-\phi(0)=\int_{0}^{t}\phi^{\prime}(u)\;du\geq 3t/8.$
An application Cauchy-Schwarz inequality yields
$\|\Phi_{i}(tu)\|_{2}\geq u^{\top}\Phi_{i}(tu)\geq 3t/8+u^{\top}\Phi_{i}(0)\geq 3t/8-\|\Phi_{i}(0)\|_{2}.$
Setting $t=\|x\|_{2}$ and $u=x/\|x\|_{2}$, we have $\|\Phi_{i}(x)\|_{2}\geq 3\|x\|_{2}/8-\|\Phi_{i}(0)\|_{2}$. This implies that $\Phi_{i}(\cdot)$ is proper. Then, by Theorem 2.2 of Ruzhansky and Sugimoto (2015) shows that $\Phi_{i}(\cdot)$ is a diffeomorphism. Therefore, we can find a unique $x_{i}$ such that ${\Phi}_{i}\left(x_{i}\right)=x_{i+1}$, which in turn allows us to derive
$p_{{\Phi}_{i}\left(X_{i}\right)}\left(x_{i+1}\right)=p_{X_{i}}\left({\Phi}_{i}^{-1}\left(x_{i}\right)\right)\cdot\operatorname{det}\left(\frac{\partial{\Phi}_{i}^{-1}\left(x_{i+1}\right)}{\partial x_{i+1}}\right)=p_{X_{i}}\left(x_{i}\right)\cdot\operatorname{det}\left(\frac{\partial x_{i}}{\partial x_{i+1}}\right)$
Consequently, we see that: for any $i\geq 1$ ,
$p_{{\Phi}_{i}\left(X_{i}\right)}\left(x_{i+1}\right)$ $-p_{X_{i+1}}\left(x_{i+1}\right)=p_{X_{i}}\left(x_{i}\right)\operatorname{det}\left(\frac{\partial x_{i}}{\partial x_{i+1}}\right)-p_{X_{i+1}}\left(x_{i+1}\right)$ (27)
$=\int\left\{p_{X_{t}\mid X_{N}}\left(x_{i}\mid x_{N}\right)\operatorname{det}\left(\frac{\partial x_{i}}{\partial x_{i+1}}\right)-p_{X_{i+1}\mid X_{N}}\left(x_{i+1}\mid x_{N}\right)\right\}p_{X_{N}}\left(x_{N}\right)\mathrm{d}x_{N}$
$=\int\Big\{1-\underbrace{\frac{p_{X_{i+1}\mid X_{N}}\left(x_{i+1}\mid x_{N}\right)}{p_{X_{i}\mid X_{N}}\left(x_{i}\mid x_{N}\right)}\operatorname{det}\left(\frac{\partial x_{i+1}}{\partial x_{i}}\right)}_{=:\mathcal{T}\left(x_{i},x_{N}\right)}\Big\}p_{X_{i},X_{N}}\left(x_{i},x_{N}\right)\operatorname{det}\left(\frac{\partial x_{i}}{\partial x_{i+1}}\right)\mathrm{d}x_{N}$
$=\int\left\{1-\mathcal{T}\left(x_{i},x_{N}\right)\right\}p_{X_{i},X_{N}}\left(x_{i},x_{N}\right)\operatorname{det}\left(\frac{\partial x_{i}}{\partial x_{i+1}}\right)\mathrm{d}x_{N}$
Step 3: Analysis of $\mathcal{T}(x_{i},x_{N})$.
The next step is then to analyze $\mathcal{T}\left(x_{i},x_{N}\right)$. Controlling $\mathcal{T}\left(x_{i},x_{N}\right)$ : Given how $X_{i+1}$ and $X_{i}$ are generated, we observe that
$\mathcal{T}\left(x_{i+1},x_{N}\right)$ $=\frac{p_{X_{i+1}\mid X_{N}}\left(x_{i+1}\mid x_{N}\right)}{p_{X_{i}\mid X_{N}}\left(x_{i}\mid x_{N}\right)}\operatorname{det}\left(\frac{\partial x_{i+1}}{\partial x_{i}}\right)=\frac{\left(\frac{1}{1-t_{i+1}}\right)^{d}\exp\left\{\frac{-\left\|x_{i+1}-t_{i+1}x_{N}\right\|_{2}^{2}}{2\left(1-t_{i+1}\right)^{2}}\right\}}{\left(\frac{1}{1-t_{i}}\right)^{d}\exp\left\{\frac{-\left\|x_{i}-t_{i}x_{N}\right\|_{2}^{2}}{2\left(1-t_{i}\right)^{2}}\right\}}\operatorname{det}\left(\frac{\partial x_{i+1}}{\partial x_{i}}\right)$
$=\underbrace{\left(\frac{1-t_{i}}{1-t_{i+1}}\right)^{d}\operatorname{det}\left(\frac{\partial x_{i+1}}{\partial x_{i}}\right)}_{=:\mathcal{T}_{1}\left(x_{i},x_{N}\right)}\underbrace{\exp\left\{\frac{\left\|x_{i}-t_{i}x_{N}\right\|_{2}^{2}}{2\left(1-t_{i}\right)^{2}}-\frac{\left\|x_{i+1}-t_{i+1}x_{N}\right\|_{2}^{2}}{2\left(1-t_{i+1}\right)^{2}}\right\}}_{=:\mathcal{T}_{2}\left(x_{t},x_{0}\right)}$
leaving us with two terms to control. Let us first study the term $\mathcal{T}_{1}\left(x_{t},x_{0}\right)$, towards which we see that
$\mathcal{T}_{1}\left(x_{t},x_{0}\right)$ $=\left(\frac{1-t_{i}}{1-t_{i+1}}\right)^{d}\operatorname{det}\left(\frac{\partial(x_{i}+\eta_{i}\widehat{v}_{i}(x_{i}))}{\partial x_{i}}\right)$
$=\left(\frac{1-t_{i}}{1-t_{i+1}}\right)^{d}\operatorname{det}\left(I+\eta_{t}\frac{\partial}{\partial x_{t}}v_{i}\left(x_{i}\right)+\eta_{i}\left(\frac{\partial}{\partial x_{i}}\widehat{v}_{i}\left(x_{i}\right)-\frac{\partial}{\partial x_{i}}v_{i}\left(x_{i}\right)\right)\right)$
$\stackrel{{\scriptstyle(a)}}{{=}}\left(\frac{1-t_{i}}{1-t_{i+1}}\right)^{d}\operatorname{det}\left(\left(1-\frac{\eta_{i}}{1-t_{i}}\right)I+\frac{\eta_{i}t_{i}}{(1-t_{i})^{3}}\operatorname{Cov}_{1\mid t_{i}}(x)+\eta_{t}\varepsilon_{i}^{J}\left(x_{i}\right)\right)$
$=\operatorname{det}\left(\frac{1-t_{i}-\eta_{i}}{1-t_{i+1}}I+\frac{\eta_{i}t_{i}}{(1-t_{i+1})(1-t_{i})^{2}}\operatorname{Cov}_{1\mid t_{i}}\left(x_{i}\right)+\frac{\eta_{i}(1-t_{i})}{(1-t_{i+1})}\varepsilon_{i}^{J}\left(x_{t}\right)\right)$
$\stackrel{{\scriptstyle(b)}}{{=}}\operatorname{det}\left(I+\frac{\eta_{i}t_{i}}{(1-t_{i+1})(1-t_{i})^{2}}\operatorname{Cov}_{1\mid t_{i}}\left(x_{i}\right)+\frac{\eta_{i}(1-t_{i})}{(1-t_{i+1})}\varepsilon_{i}^{J}\left(x_{i}\right)\right).$
Here, (a) arises from Tweedie’s formula, whereas (b) follows since
$\frac{1-t_{i}-\eta_{i}}{1-t_{i+1}}=\frac{1-t_{i+1}}{1-t_{i+1}}=1.$
Next, we turn attention to the term $\mathcal{T}_{2}\left(x_{t},x_{0}\right)$, which satisfies
$\log\mathcal{T}_{2}\left(x_{i},x_{N}\right)$ $=\frac{\left\|x_{i}-t_{i}x_{N}\right\|_{2}^{2}}{2\left(1-t_{i}\right)^{2}}-\frac{\left\|x_{i+1}-t_{i+1}x_{N}\right\|_{2}^{2}}{2\left(1-t_{i+1}\right)^{2}}$ (28)
$=\frac{\left\|x_{i}-t_{i}x_{N}\right\|_{2}^{2}}{2\left(1-t_{i}\right)^{2}}-\frac{\left\|x_{i}+\eta_{i}\widehat{v}_{i}(x_{i})-t_{i+1}x_{N}\right\|_{2}^{2}}{2\left(1-t_{i+1}\right)^{2}}$
We focus on the term $x_{i}+\eta_{i}\widehat{v}_{i}(x_{i})-t_{i+1}x_{N}$:
$x_{i}+\eta_{i}\widehat{v}_{i}(x_{i})-t_{i+1}x_{N}$ $=\left(1-\frac{\eta_{i}}{1-t_{i}}\right)x_{i}+\frac{\eta_{i}}{1-t_{i}}\mu_{1\mid t_{i}}\left(x_{i}\right)-(t_{i}+\eta_{i})x_{N}+\eta_{i}\left(\widehat{v}_{i}\left(x_{i}\right)-v_{i}\left(x_{i}\right)\right)$
$=\left(1-\frac{\eta_{i}}{1-t_{i}}\right)\left(x_{i}-t_{i}x_{N}\right)+\frac{\eta_{i}}{1-t_{i}}\mu_{1\mid t_{i}}\left(x_{i}\right)-\frac{\eta_{i}}{1-t_{i}}t_{i}x_{N}-\eta_{i}x_{N}+\eta_{t}\left(\widehat{v}_{i}\left(x_{i}\right)-v_{i}\left(x_{i}\right)\right)$
$=\left(\frac{1-t_{i+1}}{1-t_{i}}\right)\left(x_{i}-t_{i}x_{N}\right)+\frac{\eta_{i}}{1-t_{i}}\left(\mu_{1\mid t_{i}}\left(x_{i}\right)-x_{N}\right)+\eta_{i}\varepsilon_{i}^{\rm v}\left(x_{i}\right)$
Substitution into (28) yields
$\log\mathcal{T}_{2}\left(x_{i},x_{N}\right)$ (29)
$=-\frac{\eta_{i}}{(1-t_{i})^{2}(1-t_{i+1})}(x_{i}-t_{i}x_{N})^{\top}(\mu_{1\mid t_{i}}(x_{i})-x_{N})-\frac{\eta_{i}}{(1-t_{i})(1-t_{i+1})}(x_{i}-t_{i}\mu_{1\mid t_{i}}(x_{i}))^{\top}\varepsilon_{i}^{\rm v}(x_{i})$
$-\frac{\eta_{i}t_{i+1}}{(1-t_{i+1})^{2}}\left(\mu_{1\mid t_{i}}(x_{i})-x_{N}\right)\varepsilon_{i}^{\rm v}(x_{i})-\frac{\eta_{i}^{2}\|\mu_{1\mid t_{i}}(x_{i})-x_{N}\|_{2}^{2}}{2(1-t_{i})^{2}(1-t_{i+1})^{2}}-\frac{\eta_{i}^{2}\|\varepsilon_{i}^{\rm v}(x_{i})\|_{2}^{2}}{2(1-t_{i+1})^{2}}.$
Now Define:
$\xi(x_{i},x_{N})$ $:=-\frac{\eta_{i}(x_{i}-t_{i}x_{N})^{\top}(\mu_{1\mid t_{i}}(x_{i})-x_{N})}{(1-t_{i})^{2}(1-t_{i+1})}-\frac{\eta_{i}t_{i+1}\left(\mu_{1\mid t_{i}}(x_{i})-x_{N}\right)\varepsilon_{i}^{\rm v}(x_{i})}{(1-t_{i+1})^{2}}$
$-\frac{\eta_{i}^{2}\|\mu_{1\mid t_{i}}(x_{i})-x_{N}\|_{2}^{2}}{2(1-t_{i})^{2}(1-t_{i+1})^{2}}+\frac{\eta_{i}}{(1-t_{i})^{2}(1-t_{i+1})}\left(t_{i}+\frac{\eta_{i}}{2(1-t_{i+1})}\right)\operatorname{Tr}(\text{Cov}_{1\mid t_{i}})$
It is easy to see that
$\int\xi(x_{i},x_{N})p_{X_{N}\mid X_{i}=x_{i}}(x_{N})\;dx_{N}=0.$
Thus, based on (29), we can further simplify $\log\mathcal{T}_{2}\left(x_{i},x_{N}\right)$ as follows:
$\log\mathcal{T}_{2}\left(x_{i},x_{N}\right)=$ $\xi(x_{i},x_{N})-\frac{\eta_{i}(x_{i}-t_{i}\mu_{1\mid t_{i}}(x_{i}))^{\top}\varepsilon_{i}^{\rm v}(x_{i})}{(1-t_{i})(1-t_{i+1})}-\frac{\eta_{i}^{2}\|\varepsilon_{i}^{\rm v}(x_{i})\|_{2}^{2}}{2(1-t_{i+1})^{2}}$
$-\frac{\eta_{i}}{(1-t_{i})^{2}(1-t_{i+1})}\left(t_{i}+\frac{\eta_{i}}{2(1-t_{i+1})}\right)\operatorname{Tr}(\text{Cov}_{1\mid t_{i}})$
Now, for ease of presentation, define
$W\left(x_{i}\right)$ $=\log\operatorname{det}\left(I+\frac{\eta_{i}t_{i}}{(1-t_{i+1})(1-t_{i})^{2}}\operatorname{Cov}_{1\mid t_{i}}\left(x_{i}\right)+\frac{\eta_{i}(1-t_{i})}{(1-t_{i+1})}\varepsilon_{i}^{J}\left(x_{i}\right)\right)-\frac{\eta_{i}^{2}\|\varepsilon_{i}^{\rm v}(x_{i})\|_{2}^{2}}{2(1-t_{i+1})^{2}}$
$-\frac{\eta_{i}}{(1-t_{i})^{2}(1-t_{i+1})}\left(t_{i}+\frac{\eta_{i}}{2(1-t_{i+1})}\right)\operatorname{Tr}(\text{Cov}_{1\mid t_{i}})-\frac{\eta_{i}(x_{i}-t_{i}\mu_{1\mid t_{i}}(x_{i}))^{\top}\varepsilon_{i}^{\rm v}(x_{i})}{(1-t_{i})(1-t_{i+1})}$
Then for any $x_{i}\in\mathcal{X}_{\text{data }}$, it holds that
$\int_{x_{N}}\left(1-e^{\xi\left(x_{t},x_{N}\right)}\right)e^{W\left(x_{i}\right)}p_{X_{N}\mid X_{i}}\left(x_{N}\mid x_{i}\right)\mathrm{d}x_{N}\leq-e^{W\left(x_{i}\right)}\int_{x_{N}}\xi\left(x_{i},x_{N}\right)p_{X_{N}\mid X_{i}}\left(x_{N}\mid x_{i}\right)\mathrm{d}x_{N}=0$ (30)
where the inequality results from the elementary inequality $1-e^{x}\leq-x$ for all $x\in\mathbb{R}$. Using the above quantities and (27) we get: for any $\mathcal{A}\subseteq\mathcal{X}_{\text{data }}\cap\mathcal{E}_{i}$,
$\mathbb{P}_{\Phi_{i}\left(X_{i}\right)}(\mathcal{A})-\mathbb{P}_{X_{i+1}}(\mathcal{A})$
$=\int_{\mathcal{A}}\left\{1-\mathcal{T}\left(x_{i},x_{N}\right)\right\}p_{X_{i},X_{N}}\left(x_{i},x_{N}\right)\operatorname{det}\left(\frac{\partial x_{i}}{\partial x_{i+1}}\right)\mathrm{d}x_{N}\;{\mathrm{d}x}_{i+1}$
$=\int_{\mathcal{A}\times\mathcal{X}_{\text{data }}}\left\{1-\mathcal{T}\left(x_{i},x_{N}\right)\right\}p_{X_{N}\mid X_{i}=x_{i}}\left(x_{N}\right)p_{X_{i}}(x_{i})\operatorname{det}\left(\frac{\partial x_{i}}{\partial x_{i+1}}\right)\mathrm{d}x_{N}\;{\mathrm{d}x}_{i+1}$
$=\int_{\Phi_{i}^{-1}(\mathcal{A})\times\mathcal{X}_{\text{data }}}\left\{1-\mathcal{T}\left(x_{i},x_{N}\right)\right\}p_{X_{N}\mid X_{i}=x_{i}}\left(x_{N}\right)p_{X_{i}}(x_{i})\mathrm{d}x_{N}\mathrm{~d}x_{i}$
$=\int_{\Phi_{i}^{-1}(\mathcal{A})\times\mathcal{X}_{\text{data }}}\left\{1-e^{\xi\left(x_{i},x_{N}\right)}\cdot e^{W\left(x_{i}\right)}\right\}p_{X_{N}\mid X_{i}=x_{i}}\left(x_{N}\right)p_{X_{i}}(x_{i})\mathrm{d}x_{N}\mathrm{~d}x_{i}$
$=\int_{\Phi_{i}^{-1}(\mathcal{A})\times\mathcal{X}_{\text{data }}}\left\{\left(1-e^{\xi\left(x_{i},x_{N}\right)}\right)e^{W\left(x_{i}\right)}+\left(1-e^{W\left(x_{i}\right)}\right)\right\}p_{X_{N}\mid X_{i}=x_{i}}\left(x_{N}\right)p_{X_{i}}(x_{i})\mathrm{d}x_{N}\mathrm{~d}x_{i}$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\int_{x_{i},x_{N}\in\Phi_{i}^{-1}(\mathcal{A})\times\mathcal{X}_{\rm data}}\left\{1-e^{W\left(x_{i}\right)}\right\}p_{X_{i},X_{N}}\left(x_{i},x_{N}\right)\mathrm{d}x_{i}\;{\mathrm{d}x}_{N}$
$\leq\int_{x_{i}\in\Phi_{i}^{-1}(\mathcal{A})}-W\left(x_{i}\right)p_{X_{i}}\left(x_{i}\right)\mathrm{d}x_{i},$
where, ($a$) invokes (30), and the last inequality follows from the elementary inequality $1-e^{x}\leq-x$.
Step 4: Analyzing $W(x_{i})$.
First, we introduce the following lemma from (Liang et al., 2025).
Lemma B.1 (Lemma 5, Liang et al. (2025)).
Let $A\in\mathbb{R}^{d\times d}$ be a positive-definite matrix, and $\Delta\in\mathbb{R}^{d\times d}$ be any square matrix. Suppose $\eta\|\Delta\|_{2}\leq 1/4$, where $\eta\in(0,1)$. Then it holds that
$\log\det(I+\eta A+\eta\Delta)\geq\eta(\operatorname{Tr}(A)+\operatorname{Tr}(\Delta))-4\eta^{2}(\|A\|_{F}^{2}+\|\Delta\|_{F}^{2}).$
Using the above lemma we have
$-\log\operatorname{det}\left(I+\frac{\eta_{i}t_{i}}{(1-t_{i+1})(1-t_{i})^{2}}\operatorname{Cov}_{1\mid t_{i}}\left(x_{i}\right)+\frac{\eta_{i}(1-t_{i})}{(1-t_{i+1})}\varepsilon_{i}^{J}\left(x_{i}\right)\right)$
$\leq\frac{4\eta_{i}^{2}t_{i}^{2}}{(1-t_{i+1})^{2}(1-t_{i})^{4}}\|\text{Cov}_{1\mid t_{i}}\|_{F}^{2}+\frac{4\eta_{i}^{2}(1-t_{i})^{2}}{(1-t_{i+1})^{2}}\|\varepsilon_{i}^{J}(x_{i})\|_{F}^{2}$
$\quad-\frac{\eta_{i}t_{i}}{(1-t_{i+1})(1-t_{i})^{2}}\operatorname{Tr}(\text{Cov}_{1\mid t_{i}})-\frac{\eta_{i}(1-t_{i})}{(1-t_{i+1})}\operatorname{Tr}(\varepsilon_{i}^{J}(x_{i})),$
provided that the inequality
$\frac{\eta_{i}(1-t_{i})}{(1-t_{i+1})}\|\varepsilon_{i}^{J}(x_{i})\|_{2}\leq\frac{1}{4}$
holds for all $i$. This is indeed the case in our setting, due to Assumption 4.4(a), and
$\frac{1-t_{i}}{1-t_{i+1}}=\frac{1}{1-\frac{\eta_{i}}{1-t_{i}}}\leq 2$
due to the inequality $\frac{\eta_{i}}{1-t_{i}}\leq 2h\leq 1/2$, where $h$ is defined in (15). This yields
$-W\left(x_{i}\right)$ $=-\log\operatorname{det}\left(I+\frac{\eta_{i}t_{i}}{(1-t_{i+1})(1-t_{i})^{2}}\operatorname{Cov}_{1\mid t_{i}}\left(x_{i}\right)+\frac{\eta_{i}(1-t_{i})}{(1-t_{i+1})}\varepsilon_{i}^{J}\left(x_{i}\right)\right)+\frac{\eta_{i}^{2}\|\varepsilon_{i}^{\rm v}(x_{i})\|_{2}^{2}}{2(1-t_{i+1})^{2}}$ (31)
$+\frac{\eta_{i}}{(1-t_{i})^{2}(1-t_{i+1})}\left(t_{i}+\frac{\eta_{i}}{2(1-t_{i+1})}\right)\operatorname{Tr}(\text{Cov}_{1\mid t_{i}})+\frac{\eta_{i}(x_{i}-t_{i}\mu_{1\mid t_{i}}(x_{i}))^{\top}\varepsilon_{i}^{\rm v}(x_{i})}{(1-t_{i})(1-t_{i+1})}$
$\leq\frac{4\eta_{i}^{2}t_{i}^{2}}{(1-t_{i+1})^{2}(1-t_{i})^{4}}\|\text{Cov}_{1\mid t_{i}}(x_{i})\|_{F}^{2}+\frac{4\eta_{i}^{2}(1-t_{i})^{2}}{(1-t_{i+1})^{2}}\|\varepsilon_{i}^{J}(x_{i})\|_{F}^{2}+\frac{\eta_{i}^{2}\|\varepsilon_{i}^{\rm v}(x_{i})\|_{2}^{2}}{2(1-t_{i+1})^{2}}$
$\quad+\frac{\eta_{i}(x_{i}-t_{i}\mu_{1\mid t_{i}}(x_{i}))^{\top}\varepsilon_{i}^{\rm v}(x_{i})}{(1-t_{i})(1-t_{i+1})}+\frac{\eta_{i}}{(1-t_{i})^{2}(1-t_{i+1})}\left(t_{i}+\frac{\eta_{i}}{2(1-t_{i+1})}-t_{i}\right)\operatorname{Tr}(\text{Cov}_{1\mid t_{i}}(x_{i}))$
$\quad-\frac{\eta_{i}^{2}(1-t_{i})}{(1-t_{i+1})}\operatorname{Tr}(\varepsilon_{i}^{J}(x_{i}))$
$\leq\frac{4\eta_{i}^{2}t_{i}^{2}}{(1-t_{i+1})^{2}(1-t_{i})^{4}}\|\text{Cov}_{1\mid t_{i}}(x_{i})\|_{F}^{2}+\frac{4\eta_{i}^{2}(1-t_{i})^{2}}{(1-t_{i+1})^{2}}\|\varepsilon_{i}^{J}(x_{i})\|_{F}^{2}+\frac{\eta_{i}^{2}\|\varepsilon_{i}^{\rm v}(x_{i})\|_{2}^{2}}{2(1-t_{i+1})^{2}}$
$\quad+\frac{\eta_{i}^{2}}{2(1-t_{i})^{2}(1-t_{i+1})^{2}}\operatorname{Tr}(\text{Cov}_{1\mid t_{i}}(x_{i}))+\underbrace{\frac{\eta_{i}(x_{i}-t_{i}\mu_{1\mid t_{i}}(x_{i}))^{\top}\varepsilon_{i}^{\rm v}(x_{i})}{(1-t_{i})(1-t_{i+1})}}_{\Delta(\varepsilon_{i}^{\rm v}(x_{i}))}$
$\quad-\frac{\eta_{i}^{2}(1-t_{i})}{(1-t_{i+1})}\operatorname{Tr}(\varepsilon_{i}^{J}(x_{i}))$
$\leq\frac{4\eta_{i}^{2}t_{i}^{2}}{(1-t_{i+1})^{2}(1-t_{i})^{4}}\|\text{Cov}_{1\mid t_{i}}(x_{i})\|_{F}^{2}+\frac{4\eta_{i}^{2}(1-t_{i})^{2}}{(1-t_{i+1})^{2}}\|\varepsilon_{i}^{J}(x_{i})\|_{F}^{2}+\frac{\eta_{i}^{2}\|\varepsilon_{i}^{\rm v}(x_{i})\|_{2}^{2}}{2(1-t_{i+1})^{2}}$
$\quad+\frac{\eta_{i}^{2}}{2(1-t_{i})^{2}(1-t_{i+1})^{2}}\operatorname{Tr}(\text{Cov}_{1\mid t_{i}}(x_{i}))+\underbrace{\frac{\eta_{i}(1-t_{i})}{(1-t_{i+1})}\frac{(x_{i}-t_{i}\mu_{1\mid t_{i}}(x_{i}))^{\top}\varepsilon_{i}^{\rm v}(x_{i})}{(1-t_{i})^{2}}}_{\Delta(\varepsilon_{i}^{\rm v}(x_{i}))}$
$\quad-\frac{\eta_{i}^{2}(1-t_{i})}{(1-t_{i+1})}\operatorname{Tr}(\varepsilon_{i}^{J}(x_{i}))$
Step 5: Controlling the error arising from $\Delta(\varepsilon_{i}^{\rm v}(x_{i}))$.
We start with the following integral quantity for any set $\mathcal{A}$:
$\int_{x_{i}\in\mathcal{A}}\frac{1}{(1-t_{i})^{2}}(x_{i}-t_{i}\mu_{1\mid t_{i}}(x_{i}))^{\top}\varepsilon^{\rm v}_{i}(x_{i})p_{X_{i}}(x_{i})\;{\mathrm{d}x}_{i}$ (32)
$=\int_{x_{i},x_{N}\in\mathcal{A}\times\mathcal{X}_{\rm data}}\frac{1}{(1-t_{i})^{2}}(x_{i}-t_{i}\mu_{1\mid t_{i}}(x_{i}))^{\top}\varepsilon^{\rm v}_{i}(x_{i})p_{X_{i},X_{N}}(x_{i},x_{N})\;{\mathrm{d}x}_{i}\;{\mathrm{d}x}_{N}$
$=\int_{x_{i},x_{N}\in\mathcal{A}\times\mathcal{X}_{\rm data}}\frac{1}{(1-t_{i})^{2}}(x_{i}-t_{i}\mu_{1\mid t_{i}}(x_{i}))^{\top}\varepsilon^{\rm v}_{i}(x_{i})p_{X_{i}\mid X_{N}=x_{N}}(x_{i})p_{X_{N}}(x_{N})\;{\mathrm{d}x}_{i}\;{\mathrm{d}x}_{N}$
$=\int_{x_{i},x_{N}\in\mathbb{R}^{d}\times\mathcal{X}_{\rm data}}\frac{1}{(1-t_{i})^{2}}\left\{(x_{i}-t_{i}x_{N})^{\top}\varepsilon^{\rm v}_{i}(x_{i})p_{X_{i}\mid X_{N}=x_{N}}(x_{i})\;{\mathrm{d}x}_{i}\right\}p_{X_{N}}(x_{N})\;{\mathrm{d}x}_{N}$
$\leq\int_{x_{i},x_{N}\in\mathbb{R}^{d}\times\mathcal{X}_{\rm data}}\frac{1}{(1-t_{i})^{2}}\left\{\left|(x_{i}-t_{i}x_{N})^{\top}\varepsilon^{\rm v}_{i}(x_{i})\right|p_{X_{i}\mid X_{N}=x_{N}}(x_{i})\;{\mathrm{d}x}_{i}\right\}p_{X_{N}}(x_{N})\;{\mathrm{d}x}_{N}$
$\leq\left(\int_{x_{i},x_{N}\in\mathbb{R}^{d}\times\mathcal{X}_{\rm data}}\left\{\frac{1}{(1-t_{i})^{4}}\left|(x_{i}-t_{i}x_{N})^{\top}\varepsilon^{\rm v}_{i}(x_{i})\right|^{2}p_{X_{i}\mid X_{N}=x_{N}}(x_{i})\;{\mathrm{d}x}_{i}\right\}p_{X_{N}}(x_{N})\;{\mathrm{d}x}_{N}\right)^{1/2}$
$=\left(\int_{x_{i},x_{N}\in\mathbb{R}^{d}\times\mathcal{X}_{\rm data}}\left\{\frac{1}{(1-t_{i})^{4}}\left\langle(x_{i}-t_{i}x_{N})(x_{i}-t_{i}x_{N})^{\top},\varepsilon^{\rm v}_{i}(x_{i})\varepsilon^{\rm v}_{i}(x_{i})^{\top}\right\rangle p_{X_{i}\mid X_{N}=x_{N}}(x_{i})\;{\mathrm{d}x}_{i}\right\}p_{X_{N}}(x_{N})\;{\mathrm{d}x}_{N}\right)^{1/2}$
Now, recall that $X_{i}\mid X_{N}=x_{N}\sim N(t_{i}x_{N},(1-t_{i})^{2}I)$. Using standard techniques from Liang et al. (2025, Section B.5), we can obtain bound on (32):
$\int_{x_{i}\in\mathcal{A}}\frac{1}{(1-t_{i})^{2}}(x_{i}-t_{i}\mu_{1\mid t_{i}}(x_{i}))^{\top}\varepsilon^{\rm v}_{i}(x_{i})p_{X_{i}}(x_{i})\;{\mathrm{d}x}_{i}\leq\frac{2\varepsilon^{\rm v}_{i}}{(1-t_{i})}+\varepsilon^{\rm J,1}_{i}+\varepsilon^{\rm J,2}_{i}+\varepsilon^{\rm H}_{i}.$
Substituting this in (31), we have
$\int_{x_{i}\in\mathcal{A}}\left\{\Delta(\varepsilon^{\rm v}_{i}(x_{i}))-\frac{\eta_{i}^{2}(1-t_{i})}{(1-t_{i+1})}\operatorname{Tr}(\varepsilon_{i}^{J}(x_{i}))\right\}p_{X_{i}}(x_{i})$
$\leq\frac{2\varepsilon^{\rm v}_{i}}{(1-t_{i})}+\varepsilon^{\rm J,1}_{i}+\varepsilon^{\rm J,2}_{i}+\varepsilon^{\rm H}_{i}+\frac{\eta_{i}^{2}(1-t_{i})}{(1-t_{i+1})}\left({\int\operatorname{Tr}(\varepsilon_{i}^{J}(x_{i}))^{2}p_{X_{i}}(x_{i}){\mathrm{d}x}_{i}}\right)^{1/2}$
$\leq\frac{2\eta_{i}}{(1-t_{i+1})}(\varepsilon^{\rm v}_{i}+\varepsilon^{\rm J,1}_{i}+\varepsilon^{\rm J,2}_{i}+\varepsilon^{\rm H}_{i}).$
Here, we used the fact that $\frac{\eta_{i}^{2}(1-t_{i})}{(1-t_{i+1})}\leq\frac{\eta_{i}}{(1-t_{i+1})}$.
Step 6: Bounding the terms involving $\text{Cov}_{1\mid t_{i}}(x_{i})$.
Here we always assume that $k\geq\log d$. [^1] If $k<\log d$, then we can redefine $k$ as $\log d$. This does not change the results in Theorem 4.5 and Theorem 4.7. We begin defining some quantities first.
  • Let $\{x_{i}^{\ast}\}_{1\leq i\leq N_{\epsilon_{0}}}$ be an $\epsilon_{0}$-net of $\mathcal{X}_{\mathrm{data}}$, with $N_{\epsilon_{0}}$ denoting its cardinality. Let $\{B_{i}\}_{1\leq i\leq N_{\epsilon_{0}}}$ be a disjoint $\epsilon_{0}$-cover for $\mathcal{X}_{\mathrm{data}}$ such that $x_{i}^{\ast}\in B_{i}$ for each $i$.
  • Define the following two sets:
    $\mathcal{I}:=\left\{1\leq i\leq N_{\epsilon_{0}}:\mathbb{P}(X_{0}\in B_{i})\geq\exp(-C_{1}k\log T)\right\}$ (33)
    and
    $\mathcal{G}:=\left\{w\in\mathbb{R}^{d}:\|w\|_{2}\leq 2\sqrt{d}+\sqrt{C_{1}k\log T},\;\left|(x_{i}^{\ast}-x_{j}^{\ast})^{\top}w\right|\leq\sqrt{C_{1}k\log T}\,\|x_{i}^{\ast}-x_{j}^{\ast}\|_{2},\;\forall\,1\leq i,j\leq N_{\varepsilon_{0}}\right\}.$ (34)
    for some sufficiently large universal constant $C_{1}>0$. As it turns out, $\bigcup_{i\in\mathcal{I}}B_{i}$ and $\mathcal{G}$ form certain high-probability sets related to the random vector $X_{0}\sim p_{\mathrm{data}}$ and the standard Gaussian random vector in $\mathbb{R}^{d}$, respectively.
  • Recall (24), we can express
    $X_{t_{i}}\equiv X_{i}=t_{i}X_{N}+(1-t_{i})X_{0},$ (35)
    for some random vector $X_{0}\sim\mathcal{N}(0,I_{d})$.
Next, we define the set $\mathcal{T}_{t}:=\left\{tv_{1}+(1-t)\omega:v_{1}\in\cup_{i\in\mathcal{I}}\mathcal{B}_{i},\omega\in\mathcal{G}\right\}$.
Borrowing standard techniques and results from the proofs of Liang et al. (2025, Lemma 1) and Liang et al. (2025, Corollary 1), we have the following results
${\mathbb{P}}\left({X_{i}\notin\mathcal{T}_{t_{i}}}\right)\leq\exp\left({-C_{2}k\log N}\right),\quad\text{for $C_{2}\gg C_{\rm cover}$, and $0\leq t_{i}<1$.}$ (36)
$\operatorname{Tr}(\text{Cov}_{1\mid t_{i}}(x_{i}))\leq\frac{C_{3}(1-t_{i})^{2}}{t_{i}^{2}}k\log N,\quad\text{for all $t_{i}>0$ and $x_{i}\in\mathcal{T}_{t_{i}}$.}$ (37)
Step 7: Bound on the term involving squared Frobenius norm using linearization:
One can use (37) directly to bound the term $\mathbb{E}_{X_{i}}\left(\|\text{Cov}_{1\mid t_{i}}(X_{i})\|_{F}^{2}\right)$ in (31). However, this will lead to a quadratic dependence on $k$. Therefore, we present the following lemma that will yield a linear dependence on $k$. The proof is deferred to Appendix D.2.
Proposition B.2.
For all $i\geq 1$, the following holds:
$\frac{\eta_{i}t_{i}}{(1-t_{i+1})^{2}(1-t_{i})^{2}}\mathbb{E}(\|\text{Cov}_{1\mid t_{i}}(X_{i})\|_{F}^{2})\leq\frac{3}{4}\left(\mathbb{E}[\operatorname{Tr}(\text{Cov}_{1\mid t_{i}}(X_{i}))]-\mathbb{E}[\operatorname{Tr}(\text{Cov}_{1\mid t_{i+1}}(X_{i+1}))]\right)+\frac{6h\eta_{i}}{(1-t_{i+1})^{2}N^{10}}.$
The
Using Proposition B.2 and (37), we have
$\mathbb{P}_{\Phi_{i}\left(X_{i}\right)}(\mathcal{A})-\mathbb{P}_{X_{i+1}}(\mathcal{A})$
$\leq\frac{3\eta_{i}t_{i}}{(1-t_{i})^{2}}\left(\frac{(1-t_{i})^{2}}{t_{i}^{2}}-\frac{(1-t_{i+1})^{2}}{t_{i+1}^{2}}\right)k\log N+\frac{\eta_{i}^{2}}{2(1-t_{i+1})^{2}t_{i}^{2}}k\log N+\frac{4\eta_{i}^{2}(1-t_{i})^{2}}{(1-t_{i+1})^{2}}(\varepsilon^{\rm J,1}_{i})^{2}+\frac{\eta_{i}^{2}}{2(1-t_{i})^{2}}(\varepsilon^{\rm v}_{i})^{2}$
$\quad+\frac{2\eta_{i}}{(1-t_{i+1})}(\varepsilon^{\rm v}_{i}+\varepsilon^{\rm J,1}_{i}+\varepsilon^{\rm J,2}_{i}+\varepsilon^{\rm H}_{i})+{\mathbb{P}}(X_{i}\notin\mathcal{E}_{i})+\frac{24h\eta_{i}^{2}t_{i}}{(1-t_{i+1})^{2}(1-t_{i})^{2}N^{10}}.$
First, Lemma D.3 yields that
$\frac{(1-t_{i})^{2}}{t_{i}^{2}}-\frac{(1-t_{i+1})^{2}}{t_{i+1}^{2}}\leq\frac{2\eta_{i}}{t_{i}^{3}}.$ (38)
Using the above inequality along with (38) yields
$\mathbb{P}_{\Phi_{i}\left(X_{i}\right)}(\mathcal{A})-\mathbb{P}_{X_{i+1}}(\mathcal{A})$
$\leq\frac{6C_{3}\eta_{i}^{2}}{(1-t_{i+1})^{2}t_{i}^{2}}k\log N+\frac{C_{3}\eta_{i}^{2}}{2(1-t_{i+1})^{2}t_{i}^{2}}k\log N+\frac{C\eta_{i}^{2}(1-t_{i})^{2}}{(1-t_{i+1})^{2}}(\varepsilon^{\rm J,1}_{i})^{2}+\frac{\eta_{i}^{2}}{2(1-t_{i})^{2}}(\varepsilon^{\rm v}_{i})^{2}$
$\quad+\frac{2\eta_{i}}{(1-t_{i+1})}(\varepsilon^{\rm v}_{i}+\varepsilon^{\rm J,1}_{i}+\varepsilon^{\rm J,2}_{i}+\varepsilon^{\rm H}_{i})+\frac{24h\eta_{i}^{2}t_{i}}{(1-t_{i+1})^{2}(1-t_{i})^{2}N^{10}}.$
Recalling properties of time scheduling (15) (see Lemma D.2):
  1. 1.
    For all $i>1$, we have $\frac{\eta_{i}}{1-t_{i}}\leq h$ and $\frac{\eta_{i}}{1-t_{i+1}}\leq h$.
  2. 2.
    For all $i>1$, we have $\frac{\eta_{i}^{2}(1-t_{i})^{2}}{(1-t_{i+1})^{2}}=\frac{(1-t_{i})^{2}}{(\frac{1-t_{i}}{\eta_{i}}-1)^{2}}\leq h^{2}$, as $\frac{\eta_{i}}{(1-t_{i})}\leq h$ and $(1-t_{i})\leq 1$.
  3. 3.
    $\sum_{i:t_{i}>0}\frac{\eta_{i}^{2}}{(1-t_{i+1})^{2}t_{i}^{2}}=\sum_{i:0<t_{i}<1/2}\frac{\eta_{i}^{2}}{(1-t_{i+1})^{2}t_{i}^{2}}+\sum_{i:t_{i}\geq 1/2}\frac{\eta_{i}^{2}}{(1-t_{i+1})^{2}t_{i}^{2}}\leq 4h^{2}N$.
Putting all of these together and solving (25), we get
${\sf TV}(p_{X_{N-1}},p_{Y_{N-1}})$
$\leq C_{1}^{\prime}h^{2}Nk\log N++C_{2}^{\prime}h^{2}N\varepsilon_{\rm J,1}^{2}+C_{3}^{\prime}h^{2}N\varepsilon_{\rm v}^{2}$
$\quad+C_{4}^{\prime}h\times\sum_{i=1}^{N-1}(\varepsilon^{\rm v}_{i}+\varepsilon^{\rm J,1}_{i}+\varepsilon^{\rm J,2}_{i}+\varepsilon^{\rm H}_{i})+C_{5}^{\prime}\frac{h^{2}}{\delta^{2}N^{10}}+{\sf TV}(p_{X_{1}},p_{Y_{1}})$
$\leq C_{1}^{\prime}h^{2}Nk\log N+C_{2}^{\prime}h^{2}N\varepsilon_{\rm J,1}^{2}+C_{3}^{\prime}h^{2}N\varepsilon_{\rm v}^{2}$
$\quad+C_{4}^{\prime}hN(\varepsilon_{\rm v}+\varepsilon_{\rm J,1}+\varepsilon_{\rm J,2}+\varepsilon_{\rm H})+C_{5}^{\prime}\frac{h^{2}}{\delta^{2}N^{10}}+{\sf TV}(p_{X_{1}},p_{Y_{1}}),$
for sufficiently large $N$.
Handling ${\sf TV}(p_{X_{1}},p_{Y_{1}})$ separately:
Recall that (31) is also valid for $t_{i+1}=\delta,t_{i}=0$. Therefore, we have
$-W(x_{0})$
$\leq\frac{4\delta^{2}}{(1-\delta)^{2}}\|\varepsilon_{0}^{J}(x_{0})\|_{F}^{2}+\frac{\delta^{2}\|\varepsilon_{0}^{\rm v}(x_{0})\|_{2}^{2}}{2(1-\delta)^{2}}+\frac{\delta^{2}}{2(1-\delta)^{2}}{\sf Trace}(\text{Cov}_{1})+\underbrace{\frac{\delta}{1-\delta}x_{0}^{\top}\varepsilon_{0}^{\rm v}(x_{0})}_{\Delta(\varepsilon_{0}^{\rm v}(x_{0}))}$
$\quad-\frac{\delta^{2}}{(1-\delta)}{\sf Trace}(\varepsilon_{0}^{J}(x_{0})).$
Therefore, applying similar argument, we get
${\sf TV}(p_{X_{1}},p_{Y_{1}})\leq A_{1}\delta^{2}{\sf Trace}(\text{Cov}_{1})+A_{2}\delta^{2}(\varepsilon^{\rm v}_{0})^{2}+A_{3}\delta^{2}(\varepsilon^{\rm J,1}_{0})^{2}+A_{4}(\varepsilon^{\rm v}_{0}+\varepsilon^{\rm J,1}_{0}+\varepsilon^{\rm J,2}_{0}+\varepsilon^{\rm H}_{0})+\underbrace{{\sf TV}(p_{X_{0}},p_{Y_{0}})}_{=0}.$
Now we recall that $\delta<1/N$. Therefore, combining all of the above inequalities and (16) gives
${\sf TV}(p_{X_{N-1}},p_{Y_{N-1}})$
$\leq C_{1}^{\prime\prime}\frac{k\log^{3}(1/\delta)}{N}+C_{2}^{\prime\prime}\frac{\log^{2}(1/\delta)}{N}\varepsilon_{\rm J,1}^{2}+C_{3}^{\prime\prime}\frac{\log^{2}(1/\delta)}{N}\varepsilon_{\rm v}^{2}$
$\quad+C_{4}^{\prime\prime}\log(1/\delta)(\varepsilon_{\rm v}+\varepsilon_{\rm J,1}+\varepsilon_{\rm J,2}+\varepsilon_{\rm H})+C_{5}^{\prime\prime}\delta^{-2}N^{-10}$
$\quad+A_{1}\delta^{2}{\operatorname{Tr}(\text{Cov}_{1})}+A_{2}\delta^{2}(\varepsilon^{\rm v}_{0})^{2}+A_{3}\delta^{2}(\varepsilon^{\rm J,1}_{0})^{2}+A_{4}(\varepsilon^{\rm v}_{0}+\varepsilon^{\rm J,1}_{0}+\varepsilon^{\rm J,2}_{0}+\varepsilon^{\rm H}_{0}),$
where $\text{Cov}_{1}:=\text{Cov}_{X_{1}\sim p_{1}}(X_{1})$. Now, note that one can estimate $v_{0}(x)=\mu_{p_{X_{1}}}-x$ by $\widehat{v}_{0}(x):=\widehat{\mu}_{1}-x$, where $\widehat{\mu}_{1}=\sum_{i=1}^{n}x_{i,1}/n$ and $n$ is the (training) sample size. In this case, we have
  • $\varepsilon^{\rm J,1}_{0}=\varepsilon^{\rm J,2}_{0}=\varepsilon^{\rm H}=0$.
  • $(\varepsilon^{\rm v}_{0})^{2}=\mathbb{E}\|\widehat{\mu}_{1}-\mu_{p_{X_{1}}}\|_{2}^{2}=O\left(\frac{{\sf Trace}(\text{Cov}_{1})}{n}\right)$.
This finishes the proof of Theorem 4.5.
B.3 Proof of convergence of STOC-RF (Theorem 4.7)
The proof of this result essentially follows by reduction and leverages key results of (Liang et al., 2025). We start by recalling some key facts: $t(\tau)=\frac{\sqrt{\omega_{\tau}}}{\sqrt{1-\omega_{\tau}}+\sqrt{\omega_{\tau}}}$, and $\frac{Y^{\prime}_{\tau}}{\sqrt{\omega_{\tau}}}\overset{d}{=}\frac{\tilde{Z}_{t(\tau)}}{t(\tau)}$. With this time transformation, we have
$\alpha_{\tau}=\frac{\omega_{\tau}}{\omega_{\tau-1}}=\frac{(t_{i}/\sigma_{t_{i}})^{2}}{(t_{i+1}/\sigma_{t_{i+1}})^{2}}=\frac{R_{i}^{2}}{R_{i+1}^{2}},\quad\text{ where $t_{i}=t(\tau)$}.$ (39)
[…]
$\hat{Y}^{\prime}_{\tau-1}=\frac{1}{\sqrt{\alpha_{\tau}}}\left\{{\hat{Y}^{\prime}_{\tau}+\delta_{\tau}\widehat{s}_{Y^{\prime}_{\tau}}(\hat{Y}^{\prime}_{\tau})+\nu_{\tau}\xi_{\tau}}\right\},\hat{Y}^{\prime}_{N}\sim N(0,I_{d})$ (40)
$\delta_{\tau}=1-\alpha_{\tau},\nu_{\tau}=\sqrt{\frac{(\alpha_{\tau}-\omega_{\tau})(1-\alpha_{t})}{1-\omega_{\tau}}},$
$\tau=N,N-1,\ldots,1,\quad\{\xi_{\tau}\}_{\tau\geq 1}\overset{i.i.d.}{\sim}N(0,I_{d}),$
Now, if we substitute all the quantities of the above sampler by their corresponding RF versions, i.e., we get $\hat{Y}_{\tau}\rightarrow Y_{t}/\sigma_{t},\widehat{s}_{Y^{\prime}_{\tau}}(y)\rightarrow\sigma_{t}\widehat{s}_{t_{i}}(Y_{t_{i}})$, and $\alpha_{\tau}\rightarrow\frac{R_{i}^{2}}{R_{i+1}^{2}}$, then it leads to the following sampler:
$\frac{Y_{t_{i+1}}}{\sigma_{t_{i+1}}}=\frac{R_{i+1}}{R_{i}}\left\{{\frac{Y_{t_{i}}}{\sigma_{t_{i}}}+\eta_{i}\sigma_{t_{i}}\widehat{s}_{t_{i}}(Y_{t_{i}})+\sqrt{\psi_{i}}W_{t_{i}}}\right\},$ (41)
$\frac{Y_{t_{0}}}{\sigma_{t_{0}}}\sim N(0,I_{d}),\quad\{W_{t_{i}}\}_{i\geq 0}\overset{i.i.d.}{\sim}N(0,I_{d}),$
$i=0,1,\ldots,N-1,$
where $\eta_{i}=1-\frac{R_{i}^{2}}{R_{i+1}^{2}},\psi_{i}=\frac{R_{i}^{2}}{R_{i+1}^{2}}.\frac{1-R_{i+1}^{2}}{1-R_{i}^{2}}.\left({1-\frac{R_{i}^{2}}{R_{i+1}^{2}}}\right)$, and $\widehat{s}_{t_{i}}(x)=\frac{t_{i}\widehat{v}_{t_{i}}(x)-x}{1-t_{i}}$. The above is exactly the sampler described in (21). In other words, we have equivalence between two above samplers. Therefore, we can use the convergence result of (40).
Defining few auxiliary processes:
We define the auxiliary processes in exact way defined in Liang et al. (2025, Section C.1).
  • First, we define $\{Y^{\star}_{\tau}\}_{\tau=N}^{1}$ by
    $Y^{\star}_{N}=Y^{\prime}_{N}\sim N(0,I_{d}),\quad Y^{\star}_{\tau-1}=\frac{1}{\sqrt{\alpha_{\tau}}}\left({Y^{\star}_{\tau}+\delta_{\tau}s_{Y^{\prime}_{\tau}}(Y^{\star}_{\tau})+\nu_{\tau}\xi_{\tau}}\right).$
    Note that, the above update rule is just the analog of $\{\hat{Y}_{\tau}\}_{\tau}$ process with the true score.
  • We construct an auxiliary process $\bar{Y}_{\tau}$ that follows the same transition dynamics as $Y_{\tau}^{\star}$:
[…]
for any $y_{\tau-1}^{-}\neq\infty$, where we recall that $a\wedge b:=\min\{a,b\}$. It is straightforward to show that
$p_{\bar{Y}_{\tau}}\left(y_{\tau}\right)=\int_{\mathbb{R}^{d}}\left(p_{Y^{\prime}_{\tau}}\left(y_{\tau}^{-}\right)\wedge p_{\bar{Y}_{\tau}^{-}}\left(y_{\tau}^{-}\right)\right)\delta\left(y_{\tau}-y_{\tau}^{-}\right)\mathrm{d}y_{\tau}^{-}=p_{Y^{\prime}_{\tau}}\left(y_{\tau}\right)\wedge p_{\bar{Y}_{\tau}^{-}}\left(y_{\tau}\right)$ (43)
for any $y_{\tau}\neq\infty$, where $\delta(\cdot)$ denotes the Dirac measure.
  • To account for the score estimation error, we introduce another auxiliary process $\widetilde{Y}_{\tau}$ based on the dynamics of $Y_{\tau}$ :
[…]
Next, we observe that for $t_{i}=t(\tau)$
$(\varepsilon^{s}_{\tau})^{2}:=\mathbb{E}\|\widehat{s}_{Y^{\prime}_{\tau}}(Y^{\prime}_{\tau})-s_{Y^{\prime}_{\tau}}(Y^{\prime}_{\tau})\|_{2}^{2}=\sigma_{t_{i}}^{2}\mathbb{E}\|\widehat{s}_{t_{i}}(X_{t_{i}})-s_{t_{i}}(X_{t_{i}})\|_{2}^{2}\leq(\varepsilon^{\rm s}_{i})^{2}.$ (45)
Now, we can exactly replicate the analysis in Liang et al. (2025, Section C.2) till Equation (107) in Step 6, which yields the following for a universal constant $L>0$:
${\sf KL}\left({p_{\bar{Y}_{1}}\|p_{\widetilde{Y}_{1}}}\right)\leq\sum_{\tau=2}^{N}L(1-\omega_{\tau})(\varepsilon_{\tau}^{s})^{2}\lesssim\frac{\log N}{N}\sum_{t_{i}=t(2)}^{t(N)}(\varepsilon^{\rm s}_{i})^{2}\lesssim\varepsilon_{\rm s}^{2}\log N.$
After this, Step 7 of the proof follows verbatim to yield the final bound in Theorem 4.7.
B.4 Description of the new RF sampler in Remark 4.8
The DDIM sampler considered in Liang et al. (2025) takes the form
$\hat{Y}^{\prime}_{\tau-1}=\frac{1}{\sqrt{\alpha_{\tau}}}\left\{{\hat{Y}^{\prime}_{\tau}+\delta_{\tau}\widehat{s}_{Y^{\prime}_{\tau}}(\hat{Y}^{\prime}_{\tau})}\right\},\;\hat{Y}^{\prime}_{N}\sim N(0,I_{d})$ (46)
$\delta_{\tau}=\frac{1-\alpha_{\tau}}{1+\sqrt{\frac{\alpha_{\tau}-\omega_{\tau}}{1-\omega_{\tau}}}},$
$\tau=N,N-1,\ldots,1.$
Using the same transformations detailed in Section B.3 and plugging them in (46), we arrive at the RF sampler
$\frac{Y_{t_{i+1}}}{\sigma_{t_{i+1}}}=\frac{R_{i+1}}{R_{i}}\left\{{\frac{Y_{t_{i}}}{\sigma_{t_{i}}}+\eta_{i}\sigma_{t_{i}}\widehat{s}_{t_{i}}(Y_{t_{i}})}\right\},$ (47)
$\frac{Y_{t_{0}}}{\sigma_{t_{0}}}\sim N(0,I_{d}),$
$i=0,1,\ldots,N-1,$
where $\eta_{i}=\frac{1-\frac{R_{i}^{2}}{R_{i+1}^{2}}}{1+\sqrt{\frac{R_{i}^{2}}{R_{i+1}^{2}}.\frac{1-R_{i+1}^{2}}{1-R_{i}^{2}}}},R_{i}=t_{i}/\sigma_{t_{i}}$.
(a) Histogram and kernel density estimation (KDE) plot of time-grid (15) showing U-shaped distribution.
(b) Histogram and KDE plot of the transformed time-grid (18) showing U-shaped distribution.
Figure 7: Histogram of non-uniform time grids.
Algebraic simplification.
Using $R_{i}=t_{i}/\sigma_{t_{i}}$, we have
$\frac{R_{i+1}}{R_{i}}=\frac{t_{i+1}}{t_{i}}\frac{\sigma_{t_{i}}}{\sigma_{t_{i+1}}},\qquad\frac{R_{i}^{2}}{R_{i+1}^{2}}.\frac{1-R_{i+1}^{2}}{1-R_{i}^{2}}=\frac{t_{i}^{2}(1-t_{i+1})^{2}}{t_{i+1}^{2}(1-t_{i})^{2}},\qquad\eta_{i}=\frac{1-\frac{t_{i}^{2}\sigma_{t_{i+1}}^{2}}{t_{i+1}^{2}\sigma_{t_{i}}^{2}}}{1+\frac{t_{i}(1-t_{i+1})}{t_{i+1}(1-t_{i})}}.$
Multiplying (47) by $\sigma_{t_{i+1}}$ yields
$Y_{t_{i+1}}=\frac{t_{i+1}}{t_{i}}\Bigl(Y_{t_{i}}+\eta_{i}\,\sigma_{t_{i}}^{2}\,\widehat{s}_{t_{i}}(Y_{t_{i}})\Bigr).$
A direct calculation using $\sigma_{t}^{2}=t^{2}+(1-t)^{2}$ shows that
$\eta_{i}\sigma_{t_{i}}^{2}=\frac{(t_{i+1}-t_{i})\bigl(1-t_{i}\bigr)}{t_{i+1}}.$
Therefore, the deterministic sampler can be written equivalently as
$Y_{t_{i+1}}=\frac{t_{i+1}}{t_{i}}\,Y_{t_{i}}+\frac{(t_{i+1}-t_{i})(1-t_{i})}{t_{i}}\widehat{s}_{t_{i}}(Y_{t_{i}})=Y_{t_{i}}+\Delta_{i}\underbrace{\left\{{\frac{Y_{t_{i}}}{t_{i}}+\frac{1-t_{i}}{t_{i}}\widehat{s}_{t_{i}}(Y_{t_{i}})}\right\}}_{=\widehat{v}_{t_{i}}(Y_{t_{i}})},$ (48)
where $\Delta_{i}=t_{i+1}-t_{i}$. This expression matches exactly the form of the RF sampler in (5). However, the time steps $\Delta_{i}$ are different from those we propose in (15), which we used to obtain the RF convergence guarantees given in Theorem 4.5. Although the time-schedule is still U-shaped as shown in Figure 7 (b), however, it is not symmetric and, as a result, might take more discretization steps at either the beginning or the end of the sampling stage. Whether a symmetric or asymmetric schedule is better remains an open problem. Furthermore, we observe that empirically it appears to be relatively sensitive to the choice of the parameters $c_{0}$ and $c_{1}$ in (18).
Appendix C Proof of Lemmas related to Stochastic Localization (Section 3)
C.1 Proof of Lemma 3.2 and Lemma 3.3
We first show that the solution to the SDE (11) takes the form
$\frac{Y^{\prime}_{\tau}}{\sqrt{\omega_{\tau}}}=X_{1}+\widetilde{B}_{\frac{1-\omega_{\tau}}{\omega_{\tau}}}^{\prime},$ (49)
where $\omega_{\tau}=\exp(-2\int_{0}^{\tau}\beta(u)du)$ and $\left\{{\widetilde{B}^{\prime}_{\tau}}\right\}_{\tau\geq 0}$ is a standard Brownian motion. First, it is easy to see that the solution of SDE (11) is
$\frac{Y^{\prime}_{\tau}}{\sqrt{\omega_{\tau}}}=X_{1}+\underbrace{\int_{0}^{\tau}\sqrt{2\beta(t)}\exp\left({\int_{0}^{t}\beta(u)du}\right)dB^{\prime}_{t}}_{:=M_{\tau}}.$ (50)
We note that $M_{\tau}$ is continuous local Martingale with $M_{0}=0$. Also, the quadratic variation of $M_{\tau}$ is
$\left\langle M\right\rangle_{\tau}=\int_{0}^{\tau}2\beta(t)\exp\left({2\int_{0}^{t}\beta(u)du}\right)dt=\exp\left({2\int_{0}^{\tau}\beta(u)du}\right)-1=\frac{1-\omega_{\tau}}{\omega_{\tau}}.$
Then, by Dambis–Dubins–Schwarz theorem (Revuz and Yor, 2013, Theorem 1.6) there exists a Brownian motion $\left\{{\widetilde{B}_{u}^{\prime}}\right\}_{u\geq 0}$ such that $\left\{{M_{\tau}}\right\}_{\tau\geq 0}=\left\{{\widetilde{B}_{\frac{1-\omega_{\tau}}{\omega_{\tau}}}^{\prime}}\right\}_{\tau\geq 0}$. Therefore, we have (49) by substituting $M_{\tau}$ in (50).
Now, we define two processes : $\hat{U}_{s}:=sY^{\prime}_{\tau(s)}/\sqrt{\omega_{\tau(s)}}=sX_{1}+s\widetilde{B}^{\prime}_{\frac{1-\omega_{\tau(s)}}{\omega_{\tau(s)}}}$, and $\tilde{U}_{s}:=s\tilde{X}_{t(s)}/t(s)=sX_{1}+s\tilde{B}_{\frac{(1-t(s))^{2}}{t^{2}(s)}}$. Now, under the prescribed time change in Lemma 3.2 and Lemma 3.3, we have
$s\widetilde{B}^{\prime}_{\frac{1-\omega_{\tau(s)}}{\omega_{\tau(s)}}}=s\widetilde{B}^{\prime}_{1/s},\quad\text{and}\quad s\tilde{B}_{\frac{(1-t(s))^{2}}{t^{2}(s)}}=s\tilde{B}_{1/s}.$
Note, that both $\left\{{s\widetilde{B}^{\prime}_{1/s}}\right\}_{s\geq 0}$ and $\left\{{s\tilde{B}_{1/s}}\right\}_{s\geq 0}$ are Brownian motions. Hence, $\left\{{\hat{U}_{s}}\right\}_{s\geq 0}$ and $\left\{{\tilde{U}_{s}}\right\}_{s\geq 0}$ are equivalent to SL process.
C.2 Proof of Lemma 3.5: Equivalence of SDE (6) and and SDE (12)
SDE (6) to SDE (12):
We begin with the solution $\tilde{Z}_{t}$ of STOC-RF SDE (6):
$d\tilde{Z}_{t}=\left\{{\frac{\tilde{Z}_{t}}{t}+2\left({\frac{1-t}{t}}\right)s_{t}(\tilde{Z}_{t})}\right\}dt+\sqrt{2\left({\frac{1-t}{t}}\right)}dB_{t},$ (51)
where $s_{t}(\cdot)$ is the score of $\tilde{Z}_{t}$, and $\left\{{B_{t}}\right\}_{t\geq 0}$ is Brownian motion.
Recall the transformation $t(\tau)=\frac{\sqrt{\omega_{\tau}}}{\sqrt{\omega_{\tau}}+\sqrt{1-\omega_{\tau}}}$, where $\omega(\tau)=\exp\left({-2\int_{0}^{\tau}\beta(u)du}\right)$.
Now, we define $\overleftarrow{Y}_{\tau}:=\frac{\sqrt{\omega_{N-\tau}}\tilde{Z}_{t(N-\tau)}}{t(N-\tau)}$. We first begin by noting that
$\frac{\sqrt{\omega_{N-\tau}}}{t(N-\tau)}=\frac{1}{\sigma_{t(N-\tau)}}$ (52)
We also have $\log\omega_{N-\tau}=-2\int_{0}^{N-\tau}\beta(u)du$. Therefore, we have
$\quad\frac{1}{\omega_{N-\tau}}.\frac{d\omega_{N-\tau}}{dt}=2\beta(N-\tau)\frac{d\tau}{dt}$ (53)
$\Rightarrow\frac{d\tau}{dt}\stackrel{{\scriptstyle(i)}}{{=}}\frac{1}{2\beta(N-\tau)\omega(N-\tau)}.\frac{d\omega_{N-\tau}}{dt}=\frac{1}{2\beta(N-\tau)\omega(N-\tau)}.\frac{d(t^{2}/\sigma_{t}^{2})}{dt}$
$\Rightarrow\frac{d\tau}{dt}=\frac{1}{\beta(N-\tau)}.\frac{1-t}{t\sigma_{t}^{2}}$
Step (i) follows from (52). Next, we define the backward time $t_{\tau}:=t(N-\tau)$, and then we can write $\tilde{Z}_{t_{\tau}}=\sigma_{t_{\tau}}\overleftarrow{Y}_{\tau}$. This also yields that the score of $\overleftarrow{Y}_{\tau}$ satisfies $s^{\leftarrow}_{\tau}(\overleftarrow{Y}_{\tau})=\sigma_{t_{\tau}}s_{t}(\tilde{Z}_{t_{\tau}})$. Moreover, (14) yields that $\sigma_{t_{\tau}}s_{t}(\tilde{Z}_{t_{\tau}})=\nabla\log q_{N-\tau}(\overleftarrow{Y}_{\tau})$. Next, with (54) and (51) in mind, we can write
$d\overleftarrow{Y}_{\tau}$ $=\frac{1}{\sigma_{t_{\tau}}}d\tilde{Z}_{t_{\tau}}-\frac{2t-1}{\sigma_{t_{\tau}}^{3}}\tilde{Z}_{t_{\tau}}dt_{\tau}$
$=\left\{{\frac{\tilde{Z}_{t_{\tau}}}{t_{\tau}\sigma_{t_{\tau}}}+2\left({\frac{1-t_{\tau}}{t_{\tau}\sigma_{t_{\tau}}}}\right)s_{t_{\tau}}(\tilde{Z}_{t_{\tau}})-\frac{2t_{\tau}-1}{\sigma_{t_{\tau}}^{3}}\tilde{Z}_{t_{\tau}}}\right\}dt_{\tau}+\sqrt{2\left({\frac{1-t_{\tau}}{\sigma_{t_{\tau}}^{2}t_{\tau}}}\right)}dB_{t_{\tau}}$
$=\left\{{\frac{1-t_{\tau}}{t_{\tau}\sigma^{3}_{t_{\tau}}}\tilde{Z}_{t_{\tau}}+2\left({\frac{1-t_{\tau}}{t_{\tau}\sigma_{t_{\tau}}}}\right)s_{t_{\tau}}(\tilde{Z}_{t_{\tau}})}\right\}dt_{\tau}+\sqrt{2\left({\frac{1-t_{\tau}}{\sigma_{t_{\tau}}^{2}t_{\tau}}}\right)}dB_{t_{\tau}}$
$=\left\{{\frac{1-t_{\tau}}{t_{\tau}\sigma^{2}_{t_{\tau}}}\overleftarrow{Y}_{\tau}+2\left({\frac{1-t_{\tau}}{t_{\tau}\sigma^{2}_{t_{\tau}}}}\right)s^{\leftarrow}_{t_{\tau}}(\overleftarrow{Y}_{\tau})}\right\}\times\beta(N-\tau)\times\frac{t_{\tau}\sigma_{t_{\tau}}^{2}}{1-t_{\tau}}d\tau+\sqrt{2\left({\frac{1-t_{\tau}}{\sigma_{t_{\tau}}^{2}t_{\tau}}}\right)}\times\sqrt{\beta(N-\tau)\times\frac{t_{\tau}\sigma_{t_{\tau}}^{2}}{1-t_{\tau}}}d\overleftarrow{B}_{\tau}$
$=\beta(N-\tau)\left\{{\overleftarrow{Y}_{\tau}+2s_{\tau}^{\leftarrow}(\overleftarrow{Y}_{\tau})}\right\}d\tau+\sqrt{2\beta(N-\tau)}d\overleftarrow{B}_{\tau}$
$=\beta(N-\tau)\left\{{\overleftarrow{Y}_{\tau}+2\nabla\log q_{N-\tau}(\overleftarrow{Y}_{\tau})}\right\}d\tau+\sqrt{2\beta(N-\tau)}d\overleftarrow{B}_{\tau},$
where $\left\{{\overleftarrow{B}_{\tau}}\right\}_{\tau\geq 0}$ is another Brownian motion.
Initial condition check:
Recall that $\overleftarrow{Y}_{0}=\frac{\sqrt{\omega_{N}}\tilde{Z}_{t(N)}}{t(N)}=(\sqrt{\omega_{N}}+\sqrt{1-\omega_{N}})\tilde{Z}_{t(N)}$. Therefore, we have
$\text{Law}(\overleftarrow{Y}_{0})=\text{Law}\left[{(\sqrt{\omega_{N}}+\sqrt{1-\omega_{N}}).\left\{{t(N)X_{1}+(1-t(N))X_{0}}\right\}}\right]=\text{Law}(\sqrt{\omega_{N}}X_{1}+\sqrt{1-\omega_{N}}X_{0})=\text{Law}(Y^{\prime}_{N}).$
This shows that $\overleftarrow{Y}_{\tau}$ satisfies the backward DDPM SDE (12).
SDE (12) to SDE (6):
For the other direction, we let $\left\{{\overleftarrow{Y}_{\tau}}\right\}_{\tau\geq 0}$ be a solution of SDE (12). Now, to generate a solution of SDE (6), we need to define
$\tilde{Z}_{t}=\frac{t(N-\tau(t))\overleftarrow{Y}_{\tau(t)}}{\sqrt{\omega_{N-\tau(t)}}}=\frac{\overleftarrow{Y}_{\tau(t)}}{\sqrt{\omega_{N-\tau(t)}}+\sqrt{1-\omega_{N-\tau(t)}}},$
where $\tau(t)$ should satisfy the following:
$t(N-\tau(t))=$ $\quad\frac{\sqrt{\omega_{N-\tau(t)}}}{\sqrt{\omega_{N-\tau(t)}}+\sqrt{1-\omega_{N-\tau(t)}}}=t$ (54)
$\Rightarrow\omega_{N-\tau(t)}=\frac{t^{2}}{t^{2}+(1-t)^{2}}$
$\Rightarrow\exp\left({-2\int_{0}^{N-\tau(t)}\beta(u)du}\right)=\frac{t^{2}}{t^{2}+(1-t)^{2}}$
$\Rightarrow\int_{0}^{N-\tau(t)}\beta(u)du=\log\left({\sqrt{1+\frac{(1-t)^{2}}{t^{2}}}}\right).$
This shows that $\tilde{Z}_{t}=\sigma_{t}\overleftarrow{Y}_{\tau(t)}$. Additionally, (14) yields that the score of $\tilde{Z}_{t}$ is $s_{t}(\cdot)$. Now, for convenience, we now write $\tau$ to denote $\tau(t)$. Taking derivative w.r.t. $t$ on both sides of (54) yields
$\frac{d\tau}{dt}=\frac{1}{\beta(N-\tau)}.\frac{(1-t)}{t\sigma_{t}^{2}}.$
Using these facts, we finally get
$d\tilde{Z}_{t}$ $=\frac{(2t-1)}{\sigma_{t}}\overleftarrow{Y}_{\tau}+\sigma_{t}d\overleftarrow{Y}_{\tau}$
$=\frac{(2t-1)}{\sigma_{t}^{2}}\tilde{Z}_{t}+\sigma_{t}\left\{{\overleftarrow{Y}_{\tau}+2s_{\tau}^{\leftarrow}(\overleftarrow{Y}_{\tau})}\right\}\beta(N-\tau)d\tau+\sigma_{t}\sqrt{2\beta(N-\tau)}d\overleftarrow{B}_{\tau}$
$=\frac{(2t-1)}{\sigma^{2}_{t}}\tilde{Z}_{t}+\sigma_{t}\left\{{\frac{\tilde{Z}_{t}}{\sigma_{t}}+2\sigma_{t}s_{t}(\tilde{Z}_{t})}\right\}\frac{(1-t)}{t\sigma_{t}^{2}}dt+\sqrt{2\left({\frac{1-t}{t}}\right)}d\tilde{B}_{t}$
$=\left({\frac{2t-1}{\sigma_{t}^{2}}+\frac{1-t}{t\sigma_{t}^{2}}}\right)\tilde{Z}_{t}+2\left({\frac{1-t}{t}}\right)s_{t}(\tilde{Z}_{t})+\sqrt{2\left({\frac{1-t}{t}}\right)}d\tilde{B}_{t}$
$=\frac{\tilde{Z}_{t}}{t}+2\left({\frac{1-t}{t}}\right)s_{t}(\tilde{Z}_{t})+\sqrt{2\left({\frac{1-t}{t}}\right)}d\tilde{B}_{t}.$
Therefore, we $\tilde{Z}_{t}$ satisfies SDE (6).
C.3 Proof of Proposition 3.6
Recall that $\frac{\tilde{I}_{\theta(s)}}{a_{\theta(s)}}=X_{1}+W_{r^{2}_{\theta(s)}}$, where $r^{2}_{\theta(s)}=1/s$. Now, we define $\hat{U}_{s}:=\frac{\tilde{I}_{\theta(s)}}{a_{\theta(s)}}$. Then, we have $s\hat{U}_{s}=sX_{1}+sW_{1/s}$. Similar to the proofs of Lemma 3.2 and Lemma 3.3, the process $\left\{{\hat{W}_{s}}\right\}_{s\geq 0}:=\left\{{sW_{1/s}}\right\}_{s\geq 0}$ is also a Brownian motion. This finishes the proof by the description of the SL process in (9).
Appendix D Auxiliary results
D.1 Controlling posterior covariance of RF
We recall the SL process (9) to study the dynamics of the conditional covariance matrix ${\boldsymbol{\Sigma}}_{t}:=\text{Cov}_{1\mid t}(X_{t})$ where $X_{t}=tX_{1}+(1-t)X_{0}$ and $X_{0}\sim N(0,I_{d})$. Note that ${\boldsymbol{\Sigma}}_{t}$ is a random matrix. Recall the defintion of $\tilde{X}_{t}$ in (13). It is evident that $\tilde{\boldsymbol{\Sigma}}_{t}:=\text{Cov}_{1\mid t}(\tilde{X}_{t})\overset{d}{=}{\boldsymbol{\Sigma}}_{t}$. Due to Lemma 3.3, the SL process (9) is equivalent to the process (13) under time change $t(s):=\frac{\sqrt{s}}{1+\sqrt{s}}$. Then, one can substitute this in Lemma 3.1 to get the following result:
Lemma D.1 (Corollary 3.4).
for $t\in[0,1)$, we have $\frac{d}{dt}\mathbb{E}({\boldsymbol{\Sigma}}_{t})=-\frac{2t}{(1-t)^{3}}\mathbb{E}({\boldsymbol{\Sigma}}_{t}^{2})$.
Proof.
Now, we note that $\mathbf{A}_{s}$ and $\tilde{\boldsymbol{\Sigma}}_{t(s)}$ (so is ${\boldsymbol{\Sigma}}_{t(s)}$) are marginally equal in law. Therefore, by chain rule we have the following:
$\frac{d}{dt}\mathbb{E}({\boldsymbol{\Sigma}}_{t})$ $=\frac{d}{dt}\mathbb{E}(\tilde{\boldsymbol{\Sigma}}_{t})$
$=\frac{d}{dt}\mathbb{E}(\mathbf{A}_{s})$
$=\frac{2t}{(1-t)^{3}}\frac{d}{ds}\mathbb{E}(\mathbf{A}_{s})$
$=-\frac{2t}{(1-t)^{3}}\mathbb{E}(\mathbf{A}_{s}^{2})$
$=-\frac{2t}{(1-t)^{3}}\mathbb{E}({\boldsymbol{\Sigma}}_{t}^{2}).$
As mentioned in previous section, the above result allows to control the discretization error of the ODE and SDE sampler of RF. More concretely, Lemma D.1 allows us to bound the spectra of ${\boldsymbol{\Sigma}}_{t}^{2}$ in terms of the spectra of ${\boldsymbol{\Sigma}}_{t}$, thereby producing tighter bounds for convergence rate.
D.2 Proof of Proposition B.2
We recall the definitions of $(\mathbf{A}_{s})_{s\geq 0}$ and $(\tilde{\boldsymbol{\Sigma}}_{t(s)})_{s\geq 0}$ in Section 3 and Section D.1. Note that $(\mathbf{A}_{s})_{s\geq 0}$ and $(\tilde{\boldsymbol{\Sigma}}_{t(s)})_{s\geq 0}$ (see Section 3 and D.1) are also equivalent in law. This is simply due to the fact that $\tilde{\boldsymbol{\Sigma}}_{t(s)}=\text{Cov}_{1\mid t(s)}(\tilde{X}_{t(s)})=\text{Cov}_{1\mid t(s)}(s\tilde{X}_{t(s)}/t(s))$, and the distributional equivalence of the two processes $(U_{s})_{s\geq 0}$ and $(s\tilde{X}_{t(s)}/t(s))_{s\geq 0}$. It is known that the random matrix valued process $(\mathbf{A}_{s})_{s\geq 0}$ obeys the following SDE:
$d\mathbf{A}_{s}=-\mathbf{A}_{s}^{2}ds+\mathcal{M}_{s}^{(3)}dB_{s},$ (55)
where $\mathcal{M}_{s}^{(\ell)}=\mathbb{E}[(X_{1}-\mathbb{E}[X_{1}\mid U_{s}])^{\otimes\ell}\mid U_{s}]$. More details on this can be found in (Eldan, 2020, Section 4.2.1) or (Alberts et al., 2025, Section 3). Since the two processes $(\mathbf{A}_{s})_{s\geq 0}$ and $(\tilde{\boldsymbol{\Sigma}}_{t(s)})_{s\geq 0}$ are equivalent in law, SDE (55) immediately yields
$d\tilde{\boldsymbol{\Sigma}}_{t}=-\frac{2t}{(1-t)^{3}}\tilde{\boldsymbol{\Sigma}}_{t}^{2}dt+\tilde{\mathcal{M}}_{t}^{(3)}dB_{t^{2}/(1-t)^{2}},$ (56)
where $\tilde{\mathcal{M}}_{t}^{(\ell)}=\mathbb{E}[(X_{1}-\mathbb{E}[X_{1}\mid\tilde{X}_{t}])^{\otimes\ell}\mid\tilde{X}_{t}]$. Now, an application of Ito’s formula on (55) yields
$d({\sf Trace}(\mathbf{A}_{s}^{2}))=-2\left\langle\mathbf{A}_{s},\mathbf{A}_{s}^{2}\right\rangle ds+\left\langle\mathcal{M}_{s}^{(3)},\mathcal{M}_{s}^{(3)}\right\rangle ds+2\left\langle\mathbf{A}_{s},\mathcal{M}_{s}^{(3)}\right\rangle dB_{s}$
Taking expectation on both side we arrive at the following inequality:
$d({\sf Trace}[\mathbb{E}(\mathbf{A}_{s}^{2})])=-2\mathbb{E}\left[\left\langle\mathbf{A}_{s},\mathbf{A}_{s}^{2}\right\rangle\right]ds+\mathbb{E}\left[\left\langle\mathcal{M}_{s}^{(3)},\mathcal{M}_{s}^{(3)}\right\rangle\right]ds.$
Finally, using the fact that $\mathbf{A}_{s}$ and ${\boldsymbol{\Sigma}}_{t(s)}$ have the same law, chain rule yields the following:
$d({\sf Trace}[\mathbb{E}({\boldsymbol{\Sigma}}_{t}^{2})])=-\frac{4t}{(1-t)^{3}}\mathbb{E}\left[\left\langle{\boldsymbol{\Sigma}}_{t},{\boldsymbol{\Sigma}}_{t}^{2}\right\rangle\right]dt+\frac{2t}{(1-t)^{3}}\mathbb{E}\left[\left\langle\tilde{\mathcal{M}}_{t}^{(3)},\tilde{\mathcal{M}}_{t}^{(3)}\right\rangle\right]dt.$ (57)
We first focus on the first term on the right-hand side of (57). Using the symmetry of ${\boldsymbol{\Sigma}}_{t}$, we observe that
$\mathbb{E}\left[\left\langle{\boldsymbol{\Sigma}}_{t},{\boldsymbol{\Sigma}}_{t}^{2}\right\rangle\right]=\mathbb{E}\left[{\sf Trace}({\boldsymbol{\Sigma}}_{t}^{3})\right]$ (58)
$\leq\mathbb{E}\left[\|{\boldsymbol{\Sigma}}_{t}\|_{{\rm op}}\|{\boldsymbol{\Sigma}}_{t}\|^{2}_{F}\right]$
$\leq\mathbb{E}\left[{\sf Trace}\left({\boldsymbol{\Sigma}}_{t}\right)\|{\boldsymbol{\Sigma}}_{t}\|^{2}_{F}\right]$
$=\mathbb{E}\left[{\sf Trace}\left({\boldsymbol{\Sigma}}_{t}\right){\mathbbm{1}}\{X_{t}\in\mathcal{T}_{t}\}\|{\boldsymbol{\Sigma}}_{t}\|^{2}_{F}\right]+\mathbb{E}\left[{\sf Trace}\left({\boldsymbol{\Sigma}}_{t}\right){\mathbbm{1}}\{X_{t}\notin\mathcal{T}_{t}\}\|{\boldsymbol{\Sigma}}_{t}\|^{2}_{F}\right]$
$\leq C_{3}\frac{(1-t)^{2}}{t^{2}}(k\log N)\mathbb{E}\left[\|{\boldsymbol{\Sigma}}_{t}\|_{F}^{2}\right]+\mathbb{E}\left[{\sf Trace}\left({\boldsymbol{\Sigma}}_{t}\right){\mathbbm{1}}\{X_{t}\notin\mathcal{T}_{t}\}\|{\boldsymbol{\Sigma}}_{t}\|^{2}_{F}\right],$
where the last inequality follows due to (37). Regarding the last term in above display, Assumption 4.2 and (36) yields
$\mathbb{E}\left[{\sf Trace}\left({\boldsymbol{\Sigma}}_{t}\right){\mathbbm{1}}\{X_{t}\notin\mathcal{T}_{t}\}\|{\boldsymbol{\Sigma}}_{t}\|^{2}_{F}\right]\leq 8N^{3c_{R}}{\mathbb{P}}(X_{t}\notin\mathcal{T}_{t})\leq\frac{1}{N^{10}}.$
Therefore, we arrive at
$\mathbb{E}\left[\left\langle{\boldsymbol{\Sigma}}_{t},{\boldsymbol{\Sigma}}_{t}^{2}\right\rangle\right]\leq C_{3}\frac{(1-t)^{2}}{t^{2}}(k\log N)\mathbb{E}\left[\|{\boldsymbol{\Sigma}}_{t}\|_{F}^{2}\right]+\frac{1}{N^{10}}.$ (59)
Substituting this in (58) yields,
$d\mathbb{E}(\|{\boldsymbol{\Sigma}}_{t}\|_{F}^{2})$ $=-\frac{4t}{(1-t)^{3}}\mathbb{E}\left[\left\langle{\boldsymbol{\Sigma}}_{t},{\boldsymbol{\Sigma}}_{t}^{2}\right\rangle\right]dt+\frac{2t}{(1-t)^{3}}\mathbb{E}\left[\left\langle\tilde{\mathcal{M}}_{t}^{(3)},\tilde{\mathcal{M}}_{t}^{(3)}\right\rangle\right]dt$
$\geq-\frac{4t}{(1-t)^{3}}\mathbb{E}\left[\left\langle{\boldsymbol{\Sigma}}_{t},{\boldsymbol{\Sigma}}_{t}^{2}\right\rangle\right]dt$
$\geq-\frac{4C_{3}k\log N}{(1-t)t}\mathbb{E}\left[\|{\boldsymbol{\Sigma}}_{t}\|_{F}^{2}\right]dt-\frac{4t}{(1-t)^{3}}.\frac{1}{N^{10}}dt$
$\geq-\frac{4C_{3}k\log N}{(1-t_{i})t_{i-1}}\mathbb{E}\left[\|{\boldsymbol{\Sigma}}_{t}\|_{F}^{2}\right]dt-\frac{4t_{i}}{(1-t_{i})^{3}}.\frac{1}{N^{10}}dt,$
where the last line follows holds provided $t\in[t_{i-1},t_{i}]$ with $i>1$. Therefore, we can conclude that
$\exp\left\{\frac{4C_{3}kt\log N}{(1-t_{i})t_{i-1}}\right\}\mathbb{E}(\|{\boldsymbol{\Sigma}}_{t}\|_{F}^{2})-\exp\left\{\frac{4C_{3}kt_{i-1}\log N}{(1-t_{i})t_{i-1}}\right\}\mathbb{E}(\|{\boldsymbol{\Sigma}}_{t_{i-1}}\|_{F}^{2})$
$\geq-\frac{4t_{i}}{(1-t_{i})^{3}}.\frac{1}{N^{10}}\int_{t_{i-1}}^{t}\exp\left\{\frac{4C_{3}kt\log N}{(1-t_{i})t_{i-1}}\right\}dt$
$\geq-\frac{t_{i}t_{i-1}}{C_{3}k(1-t_{i})^{2}\log N}.\frac{1}{N^{10}}\left(\exp\left\{\frac{4C_{3}kt\log N}{(1-t_{i})t_{i-1}}\right\}-\exp\left\{\frac{4C_{3}kt_{i-1}\log N}{(1-t_{i})t_{i-1}}\right\}\right)$
Dividing both sides by $\exp\left\{\frac{4C_{3}kt_{i-1}\log N}{(1-t_{i})t_{i-1}}\right\}$, we obtain
$\exp\left\{\frac{4C_{3}k(t-t_{i-1})\log N}{(1-t_{i})t_{i-1}}\right\}\mathbb{E}(\|{\boldsymbol{\Sigma}}_{t}\|_{F}^{2})-\mathbb{E}(\|{\boldsymbol{\Sigma}}_{t_{i-1}}\|_{F}^{2})$
$\geq-\frac{t_{i}t_{i-1}}{C_{3}k(1-t_{i})^{2}\log N}.\frac{1}{N^{10}}\left(\exp\left\{\frac{4C_{3}k(t-t_{i-1})\log N}{(1-t_{i})t_{i-1}}\right\}-1\right)$
$\geq-\frac{1}{C_{3}k(1-t_{i})^{2}\log N}.\frac{1}{N^{10}}\left(\exp\left\{\frac{4C_{3}k(t-t_{i-1})\log N}{(1-t_{i})t_{i-1}}\right\}-1\right).$
Now, recall that $\frac{\eta_{t_{i}}}{(1-t_{i})t_{i}}\leq 2h$. Using this inequality coupled with (16) gives us
$\frac{4C_{3}k(t-t_{i-1})\log N}{(1-t_{i})t_{i-1}}\leq\frac{4C_{3}k\eta_{i-1}\log N}{(1-t_{i})t_{i-1}}\leq 8C_{3}kh\log N<1,$
provided $N$ is sufficiently large. Therefore, we have
$\mathbb{E}(\|{\boldsymbol{\Sigma}}_{t_{i-1}}\|_{F}^{2})$ $\leq\exp\left\{\frac{4C_{3}k(t-t_{i-1})\log N}{(1-t_{i})t_{i-1}}\right\}\mathbb{E}(\|{\boldsymbol{\Sigma}}_{t}\|_{F}^{2})$
$\quad+\frac{1}{C_{3}k(1-t_{i})^{2}\log N}.\frac{1}{N^{10}}\left(\exp\left\{\frac{4C_{3}k(t-t_{i-1})\log N}{(1-t_{i})t_{i-1}}\right\}-1\right)$
$\leq 3\mathbb{E}(\|{\boldsymbol{\Sigma}}_{t}\|_{F}^{2})+\frac{24h}{(1-t_{i})^{2}N^{10}}.$
The inequality follows from $e^{x}-1\leq x$ for $x\in(0,1)$, and $\frac{\eta_{i}}{(1-t_{i})t_{i}}\leq 2h$ (Lemma D.2). Substituting $t_{i-1}$ by $t_{i}$ in the above inequality, we get the following
$\mathbb{E}(\|{\boldsymbol{\Sigma}}_{t_{i}}\|_{F}^{2})\leq 3\mathbb{E}(\|{\boldsymbol{\Sigma}}_{t}\|_{F}^{2})+\frac{24h}{(1-t_{i+1})^{2}N^{10}},\quad t\in[t_{i},t_{i+1}],i\geq 1.$
Now we use Lemma D.1 to get
$\mathbb{E}[{\sf Trace}({\boldsymbol{\Sigma}}_{t_{i+1}})]-\mathbb{E}[{\sf Trace}({\boldsymbol{\Sigma}}_{t_{i}})]$ $=-\int_{t_{i}}^{t_{i+1}}\frac{2t}{(1-t)^{3}}\mathbb{E}(\|{\boldsymbol{\Sigma}}_{t}\|_{F}^{2})$
$\leq-\int_{t_{i}}^{t_{i+1}}\frac{2t}{3(1-t)^{3}}\mathbb{E}(\|{\boldsymbol{\Sigma}}_{t_{i}}\|_{F}^{2})dt+\frac{8h\eta_{i}}{(1-t_{i+1})^{2}N^{10}}$
$\leq-\frac{2\eta_{t_{i}}(t_{i+1}+t_{i})}{3(1-t_{i+1})^{2}(1-t_{i})^{2}}\mathbb{E}(\|{\boldsymbol{\Sigma}}_{t_{i}}\|_{F}^{2})+\frac{8h\eta_{i}}{(1-t_{i+1})^{2}N^{10}}$
Rearranging the terms in the above inequality and using $t_{i+1}+t_{i}>2t_{i}$, we finally get
$\frac{\eta_{t_{i}}t_{i}}{(1-t_{i+1})^{2}(1-t_{i})^{2}}\mathbb{E}(\|{\boldsymbol{\Sigma}}_{t_{i}}\|_{F}^{2})\leq\frac{3}{4}\left(\mathbb{E}[{\sf Trace}({\boldsymbol{\Sigma}}_{t_{i}})]-\mathbb{E}[{\sf Trace}({\boldsymbol{\Sigma}}_{t_{i+1}})]\right)+\frac{6h\eta_{i}}{(1-t_{i+1})^{2}N^{10}}.$ (60)
D.3 Properties of time scheduling (15)
Lemma D.2.
The time-steps (15) satisfies the following for $i\in\{1,\ldots,N-2\}$:
  1. (a)
    We have $\eta_{i}=ht_{i}{\mathbbm{1}}\left\{{t_{i}\leq 1/2}\right\}+h(1-t_{i+1}){\mathbbm{1}}\left\{{t_{i}>1/2}\right\}$.
  2. (b)
    We have $\frac{\eta_{i}}{1-t_{i}}\leq h$ and $\frac{\eta_{i}}{1-t_{i+1}}\leq h$.
  3. (c)
    We have $\frac{\eta_{i}^{2}(1-t_{i})^{2}}{(1-t_{i+1})^{2}}\leq h^{2}$, as $\frac{\eta_{i}}{(1-t_{i+1})}\leq h$ and $(1-t_{i})\leq 1$.
  4. (d)
    We have $\sum_{i:t_{i}>0}\frac{\eta_{i}^{2}}{(1-t_{i+1})^{2}t_{i}^{2}}=\sum_{i:0<t_{i}<1/2}\frac{\eta_{i}^{2}}{(1-t_{i+1})^{2}t_{i}^{2}}+\sum_{i:t_{i}\geq 1/2}\frac{\eta_{i}^{2}}{(1-t_{i+1})^{2}t_{i}^{2}}\leq 4h^{2}N$.
Proof.
Ghost line
  • Part (a): If $t_{i}<1/2$, we have $\eta_{i}=t_{i+1}-t_{i}=(1+h)t_{i}-t_{i}=ht_{i}$. For the case $t_{i}\geq 1/2$, we have
    $\eta_{i}$ $=t_{i+1}-t_{i}$
    $=(1-t_{i})-(1-t_{i+1})$
    $=h(1-t_{i+1}).$
  • Part (b): For $t_{i}<1/2$, we have $\frac{\eta_{i}}{1-t_{i}}=ht_{i}/(1-t_{i})\leq h$. We use the fact that $\frac{t}{1-t}<1$ for $t\leq 1/2$. Similarly, $\frac{\eta_{i}}{1-t_{i+1}}=\frac{ht_{i}}{1-t_{i+1}}$. Now, note that $t_{i}<1/2$ implies that $t_{i+1}\leq 1/2$. Therefore, $\frac{t_{i}}{1-t_{i+1}}\leq 1$, and it yields $\frac{\eta_{i}}{1-t_{i+1}}=\frac{ht_{i}}{1-t_{i+1}}<h$.
    For $t_{i}\geq 1/2$, we have $\frac{\eta_{i}}{1-t_{i}}=\frac{h(1-t_{i+1})}{1-t_{i}}<h$. Similarly, $\frac{\eta_{i}}{1-t_{i+1}}=\frac{h(1-t_{i+1})}{1-t_{i+1}}=h$. This yields the coveted bound.
  • Part (c): The proof immediately follows from Part (b).
  • Part (d): Fist, from Part (a), we can note that $\frac{\eta_{i}}{t_{i}}\leq h$. Also, Part (b) yields that $\frac{\eta_{i}}{1-t_{i+1}}\leq h$. Therefore,
    $\sum_{i:t_{i}>0}\frac{\eta_{i}^{2}}{(1-t_{i+1})^{2}t_{i}^{2}}$
    $=\sum_{i:0<t_{i}<1/2}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\frac{\eta_{i}^{2}}{t_{i}^{2}}}.\frac{1}{(1-t_{i+1})^{2}}+\sum_{i:t_{i}\geq 1/2}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\frac{\eta_{i}^{2}}{(1-t_{i+1})^{2}}}.\frac{1}{t_{i}^{2}}$
    $\leq\sum_{i:0<t_{i}<1/2}4h^{2}+\sum_{i:t_{i}\geq 1/2}4h^{2}$
    $\leq 4h^{2}N.$
Lemma D.3.
Let $t_{i+1}=t_{i}+\eta_{i}$ with $0<t_{i}\leq t_{i+1}\leq 1$. Then
$\frac{(1-t_{i})^{2}}{t_{i}^{2}}-\frac{(1-t_{i+1})^{2}}{t_{i+1}^{2}}\leq\frac{2\eta_{i}}{t_{i}^{3}}.$
Proof.
Define $f(t):=\frac{(1-t)^{2}}{t^{2}}$ for $t>0$. A direct calculation gives
$f^{\prime}(t)=\frac{2(t-1)}{t^{3}}.$
By the mean value theorem, there exists $\xi\in(t_{i},t_{i+1})$ such that
$f(t_{i})-f(t_{i+1})=-f^{\prime}(\xi)\,(t_{i+1}-t_{i})=\frac{2(1-\xi)}{\xi^{3}}\,\eta_{i}.$
Since $\xi\geq t_{i}$ and $\xi\leq 1$, we have $(1-\xi)/\xi^{3}\leq 1/t_{i}^{3}$, which implies the desired bound. ∎